### v2

# Intermittency and non-universality of pair dispersion in isothermal compressible turbulence

Abstract
Statistical properties of the pair dispersion of Lagrangian particles (tracers) in incompressible turbulent flows provide insights into transport and mixing. We explore the same in transonic to supersonic compressible turbulence of an isothermal ideal gas in two dimensions, driven by large-scale solenoidal and irrotational stirring forces, via direct numerical simulations. We find that the scaling exponents of the order-$p$ negative moments of the distribution of exit times – in particular, the doubling and halving times of pair separations – are nonlinear functions of $p$. Furthermore, the doubling and halving time statistics are different. The halving-time exponents are universal – they satisfy their multifractal model-based prediction, irrespective of the nature of the stirring. However, the doubling-time exponents are not. In the solenoidally-stirred flows, the doubling time exponents can be expressed solely in terms of the multifractal scaling exponents obtained from the structure functions of the solenoidal component of the velocity. Moreover, they depend strongly on the Mach number, Ma, as elongated patches of high vorticity emerge along shock fronts at high Ma. In contrast, in the irrotationally-stirred flows, the doubling-time exponents do not satisfy any known multifractal model-based relation, and are independent of Ma. Our findings are of potential relevance to astrophysical disks and molecular clouds wherein turbulent transport and mixing of gases often govern chemical kinetics and the rates of formation of stars and planetesimals.
keywords:
Compressible turbulence; Particle/Fluid Flow
1 Introduction
[…]
We begin by summarizing the evidence in support of Richardson’s law in incompressible turbulence. Theoretically, it can be understood as a modeling of the evolution of $R(t)$ by a diffusive process (Richardson, 1926; Benzi & Toschi, 2023) with a diffusion constant, $K(R)\sim R^{4/3}$, that stems (Benzi & Toschi, 2023) from Kolmogorov’s 1941 (K41) theory (Kolmogorov, 1941). An alternative interpretation of Richardson’s law is in terms of a dynamic scaling exponent, $z=2/3$, such that typical time scales, $\tau(r)$, of turbulent eddies of size $r$ scale as $\tau(r)\sim r^{z}$. It is now well established, from experiments and DNS, that K41 must be generalized to include intermittency, i.e., the $p$-th moment of the Eulerian velocity difference, $\delta u$, across a length scale $r$ (the order-$p$ velocity structure function), scales as $\left\langle\delta u^{p}(r)\right\rangle\sim r^{\zeta_{p}}$, where $\zeta_{p}$ is a nonlinear concave function of $p$ (Frisch, 1996). This also dictates that turbulence must, in fact, have an infinite number of dynamic exponents (L’vov et al., 1997; Mitra & Pandit, 2004, 2005; Ray et al., 2008, 2011; Pandit et al., 2008), suggesting the presence of intermittency corrections to Richardson’s law, which itself enjoys feeble support from experiments (Ott & Mann, 2000; Shnapp et al., 2018; Elsinga et al., 2022; Tan & Ni, 2022; Shnapp et al., 2023) and direct numerical simulations (DNS) (Salazar & Collins, 2009; Bitane et al., 2012, 2013; Bragg et al., 2016; Buaria et al., 2015; Elsinga et al., 2022) – the observed scaling range of $\left\langle R^{2}\right\rangle$ is at most a decade. A better approach to investigate the intermittency of pair dispersion in turbulence is to use exit times, defined as the time, $\tau_{\alpha}(R)$, it takes for $R$ to cross $\alpha R$ for the first time, where $\alpha$ is fixed. The choice $\alpha>1$ corresponds to the doubling times [^1] To be accurate, $\alpha=2$ means “doubling”. Nevertheless, in practice, any $\alpha>1$ is used (Boffetta & Sokolov, 2002a), $\tau_{\rm D}$, for which (Boffetta et al., 1999; Boffetta & Sokolov, 2002a, b)
$\left\langle\tau_{\rm D}^{-p}(R)\right\rangle\sim R^{-\chi^{\rm D}_{\rm p}}\,,\quad\text{where}\quad\chi^{\rm D}_{\rm p}=p-\zeta_{p}\,.$ (1)
This bridge relation between the dynamic exponents, $\chi^{\rm D}_{\rm p}$, and $\zeta_{p}$ has been obtained by using the multifractal model of turbulence (Parisi & Frisch, 1985; Frisch, 1996).
In compressible flows, in addition to doubling times, we need to define halving times, $\tau_{\rm H}(R)$, for which $\alpha<1$ and the corresponding scaling exponents are $\chi^{\rm H}_{\rm p}$. As the multifractal model does not distinguish between the doubling and halving times, $\chi^{\rm H}_{\rm p}$ are also expected naively to satisfy the same bridge relation,  (1). Although there have been many studies of the statistics of velocity increments in compressible turbulence (Boldyrev et al., 2002b, a; Schmidt et al., 2008; Konstandin et al., 2012; Wang et al., 2012a; Kritsuk et al., 2007; Wang et al., 2017; Jagannathan & Donzis, 2016), none have investigated the intermittency of tracer pair dispersion. We explore this via simulations of forced compressible turbulence of an isothermal ideal gas, governed by the Navier-Stokes equations. We restrict ourselves to two spatial dimensions in order to reach high resolutions.
Two-dimensional (2D) incompressible turbulence is different from its 3D counterpart because, in addition to the kinetic energy, the enstrophy is conserved in the limit of zero viscosity and forcing. Theory, experiments and DNS indicate that $\left\langle R^{2}(t)\right\rangle$ grows exponentially with time in the direct–enstrophy–cascade regime (Bennett, 1984; Rivera & Ecke, 2005; Jullien, 2003; LaCasce & Ohlmann, 2003). In 2D compressible turbulence, there is small-scale dissipation of kinetic energy via acoustic waves (Falkovich & Kritsuk, 2017). It has been shown that $\left\langle R^{2}(t)\right\rangle$ grows as a power–law beyond the dissipation scales in models of such flows, e.g., surface flows (Cressman & Goldburg, 2003; Li et al., 2024) and the 2D Burgers equation (De et al., 2024).
Here too, we unveil a power-law growth of $\left\langle R^{2}(t)\right\rangle$ across similar spatial scales. While, $\chi^{\rm H}_{\rm p}$ obey the multifractal model-based bridge relation and are universal, $\chi^{\rm D}_{\rm p}\neq\chi^{\rm H}_{\rm p}$, in disagreement with the multifractal model. If the external force is solenoidal: (a) $\chi^{\rm D}_{\rm p}=p-\zeta_{p}^{\rm s}$, where $\zeta_{p}^{\rm s}$ are the scaling exponents of structure functions of the solenoidal component of the velocity field; and (b) the values of $\chi^{\rm D}_{\rm p}$ (and consequently, $\zeta_{p}^{\rm s}$) depend on Ma. This indicates that vortical structures play the dominant role in pair dispersion, and we demonstrate that their fractal dimensions vary significantly with Ma, accounting for the Ma-dependence of $\chi^{\rm D}_{\rm p}$. If the force is purely irrotational, $\chi^{\rm D}_{\rm p}$ neither follows any known bridge relation, nor depends on Ma. In this case, the regions of large flow divergence affect pair dispersion.
In Section 2, we describe the model, and outline its parameters and the details of its DNS. In Section 3, we delineate our analyses of particle pair-dispersion and Lagrangian exit-times, and then provide the concluding remarks in Section 4.
2 Model and its numerical simulation
The equations (Haugen et al., 2022) governing the dynamics of a compressible isothermal ideal gas are:
$\partial_{t}\rho+\partial_{j}(\rho u_{j})$ $=0\,,$ (2a)
$\partial_{t}u_{i}+u_{j}\partial_{j}u_{i}$ $=-\frac{1}{\rho}\partial_{i}P+\frac{1}{\rho}\partial_{j}(\nu\rho\sigma_{ij}+\rho\zeta_{\rm bulk}\delta_{ij}\partial_{k}u_{k})+\mathcal{P}_{ij}(\xi)f_{j}\,,$ (2b)
where $\bm{u}({\bm{x}},t)$ is the velocity field, $\rho({\bm{x}},t)$ is the density field, $P$ is the pressure, $\nu$ is the kinematic shear viscosity, $\zeta_{\rm bulk}$ is the bulk viscosity, $\delta_{ij}$ is the Kronecker delta, $\mathcal{P}_{ij}(\xi)f_{j}$ is the external force, $\sigma_{ij}=(1/2)\left(\partial_{j}u_{i}+\partial_{i}u_{j}\right)-(1/2)\delta_{ij}\partial_{k}u_{k}$ is the traceless rate-of-strain tensor, and the equation of state is given by $P=c_{\rm s}^{2}\rho$, where $c_{\rm s}$ is the (constant) speed of sound. The quantity, ${\bm{f}}$, which we call external force or stirring, is white-in-time and its Fourier transform has compact support at $|{\bm{k}}|=k_{\rm f}=3$, where ${\bm{k}}$ is the wave vector. The Fourier transform of the projection operator, $\mathcal{P}_{ij}(\xi)$ is, $\hat{\mathcal{P}}_{ij}(\xi)\equiv\xi\delta_{ij}+(1-2\xi)k_{i}k_{j}/k^{2}$. We consider two cases, $\xi=0$ and $1$, which correspond, respectively, to an irrotational and a solenoidal external force (see Table 1). The assumption of isothermality holds when the medium is optically thin (transparent to radiation) in which case the heat generated by dissipation at shocks is rapidly radiated away. This is generally valid for molecular clouds which have low density and solar metallicity (Mac Low & Klessen, 2004).
We perform pseudospectral DNS of Eq.  (2) on a 2D doubly-periodic square domain of side $L=2\pi$ with $N^{2}=4096^{2}$ collocation points. The bulk viscosity follows the prescription outlined by Von Neumann & Richtmyer (1950): it is zero everywhere except near the shocks. We assume $\nu$ to be a constant. This is a common practice; see, e.g., Haugen et al. (2022); Federrath et al. (2011); Richard et al. (2022). After the flow reaches a statistically stationary state we seed it uniformly with $n=N^{2}$ Lagrangian particles or tracers at every grid point, such that the distance between a pair of nearest-neighbor tracers is the grid spacing. The position of the $\rm{m}$-th tracer, $\bm{X}^{\rm m}(t)$, evolves according to
$\frac{d\bm{X}^{\rm m}(t)}{dt}={\bm{V}}^{\rm m}(t)\quad\text{where}\quad{\bm{V}}^{\rm m}(t)=\bm{u}({\bm{x}},t)\delta({\bm{x}}-\bm{X}^{\rm m})\/.$ (3)
They cluster at the shocks and their number density maps $\rho({\bm{x}},t)$.
The Reynolds and turbulent Mach numbers are defined, respectively, as $\mbox{Re}\equiv u_{\rm rms}L/\nu$ and $\mbox{Ma}\equiv u_{\rm rms}/c_{\rm s}$, where $u_{\rm rms}$ is the root-mean-square velocity. Using Helmholtz decomposition, we write $\bm{u}=\bm{u}^{\rm s}+\bm{u}^{\rm c}$, where $\bm{u}^{\rm s}$ and $\bm{u}^{\rm c}$ are, respectively, the solenoidal (${\bm{\nabla}}\cdot\bm{u}^{\rm s}=0$) and irrotational (×${\bm{\nabla}}\times\bm{u}^{\rm c}=0$) components of the flow. The relative importance of $\bm{u}^{\rm s}$ and $\bm{u}^{\rm c}$ in the total kinetic energy is measured by the solenoidal ratio,
$\Psi\equiv\frac{\left\langle\rho|\bm{u}^{\rm s}|^{2}\right\rangle}{\left\langle\rho|{\bm{u}}|^{2}\right\rangle}\/,$ (4)
which is the ratio of the total kinetic energy in the solenoidal modes to the total kinetic energy. All these parameters of our DNS are summarized in Table 1. We use $L$ as our unit of length, and $T_{\rm E}$, the large-eddy turnover time, as our unit of time. Further details about the DNS are given in Appendix A.
Run $\xi$ $c_{\rm s}$ $u_{\rm rms}$ Ma Re $L_{\rm I}$ $T_{\rm E}$ $\Psi$ $\epsilon$ $\lambda$ $\mbox{Re}_{\lambda}$ $\eta$
S1 1 0.05 0.053 $1.05$ $6.7\times 10^{3}$ 1.66 32 $0.82$ $8.1\times 10^{-6}$ 0.29 306 0.0111
S2 1 0.02 0.050 $2.5$ $6.3\times 10^{3}$ 1.20 24 $0.68$ $9.6\times 10^{-6}$ 0.26 255 0.0107
C1 0 0.05 0.038 $0.75$ $4.7\times 10^{3}$ 1.04 28 $0.19$ $1.3\times 10^{-5}$ 0.16 123 0.0099
C2 0 0.02 0.040 $2.0$ $5.0\times 10^{3}$ 1.17 29 $0.34$ $8.7\times 10^{-6}$ 0.21 172 0.0109
Table 1: Simulation parameters. Runs S1 and S2 have a solenoidal external force whereas C1 and C2 have an irrotational force. The parameters are defined as follows: $c_{\rm s}$ is the speed of sound, $u_{\rm rms}$ is the root-mean-square velocity, Ma is the Mach number, $\mbox{Re}=(u_{\rm rms}L/\nu)$ is the box-size Reynolds number, $E(k)$ is the kinetic energy spectrum, $L_{\rm I}=\pi\sum_{k}[E(k)/k]\Big{/}\sum_{k}E(k)$ is the integral length scale, $T_{\rm E}=L_{\rm I}/u_{\rm rms}$ is the large-eddy turnover time, $\Psi$ is the solenoidal ratio defined in 0(4), $\epsilon=\left\langle\nu\rho\sigma_{ij}\partial_{j}u_{i}\right\rangle\Big{/}\left\langle\rho\right\rangle$ is the rate of dissipation of kinetic energy per unit mass, $\lambda=\sqrt{5\nu u_{\rm rms}^{2}\Big{/}\epsilon}$ is the Taylor microscale, $\mbox{Re}_{\lambda}=u_{\rm rms}\lambda\big{/}\nu$ is the Taylor-microscale Reynolds number, and $\eta=\left(\nu^{3}\big{/}\epsilon\right)^{1/4}$ is the dissipation scale. In all the runs, $\nu=5\times 10^{-5}$ and number of tracers is $n=4096^{2}$. Note that the length scales in this Table are not normalized by $L=2\pi$.
3 Results
Figure 1: Typical snapshots of $\nabla\cdot{\bf u}$, vorticity ${\bm{\omega}}=\hat{z}\cdot{\bm{\nabla}}\times\bm{u}$, and $\log_{10}\rho$, in the cases (a) S1, (b) S2, (c) C1 and (d) C2. In each plot, ${\bm{\nabla}}\cdot\bm{u}$ and ${\bm{\omega}}$ are normalized by their respective root-mean-square (rms) values. Shocks are clearly visible in the ${\bm{\nabla}}\cdot\bm{u}$ profiles as the blue filament–like structures. The vorticity field, $\omega$, is smoother in S1 than in S2 which contains nearly one-dimensional structures of intense vorticity bordering the shocks; in all figures, the areas of high vorticity lie in the vicinity of shocks. Similarly the spatial distribution of $\omega$ is more homogeneous in C1 than in C2. The values of $\rho$ are clearly enhanced near the shocks; for a given type of external forcing, $\rho$ tends to take much higher values at higher Mach numbers, Ma.
We perform four sets of DNSs, two for $\xi=1$, S1 and S2, and two for $\xi=0$, C1 and C2. Typically, subsonic and supersonic compressible flow can be qualitatively different. Hence we perform a set of simulations that are transonic, Runs S1 and C1, $\mbox{Ma}\simeq 1$, and another set S2 and C2 that are supersonic, $\mbox{Ma}>1$. In Fig. 1 we show pseudocolour plots of ${\bm{\nabla}}\cdot\bm{u}$, ${\bm{\omega}}={\bm{\nabla}}\times\bm{u}$, and $\log_{10}\rho$ for all the cases or runs. The shocks are clearly visible as filamentary structures with large negative values of ${\bm{\nabla}}\cdot\bm{u}$. Naively, we may expect that the presence of the shocks would not show up in pseudocolour plot of ${\bm{\omega}}$, but this not the case – the shocks can be seen as almost one-dimensional manifolds with large absolute values of the vorticity: This is because, although a shock front that is a straight line does not contribute to vorticity, curved shock fronts do. In our simulations, (a) the density fluctuations are larger for the supersonic cases compared to the transonic ones and (b) the vorticity is more uniformly distributed in the transonic cases than the supersonic ones, for a given type of external forcing, as observed in Fig. 1.
In compressible turbulence, we can define $p$-th order longitudinal structures functions of $\bm{u}$ as well as its components, $\bm{u}^{\rm s}$ and $\bm{u}^{\rm c}$:
$\delta u_{\parallel}(r)\equiv\left[\bm{u}({\bm{x}}+\bm{r})-\bm{u}({\bm{x}})\right]\cdot\frac{\bm{r}}{r}\/\quad$ $\text{and}\quad\delta u_{\parallel}^{\rm{s,c}}(r)\equiv\left[\bm{u}^{\rm{s,c}}({\bm{x}}+\bm{r})-\bm{u}^{\rm{s,c}}({\bm{x}})\right]\cdot\frac{\bm{r}}{r}\/;$ (5a)
$\text{with}\quad S_{\rm p}(r)=\left\langle|\delta u(r)_{\parallel}|^{p}\right\rangle\/,\quad$ $S_{\rm p}^{\rm s,c}(r)=\left\langle|\delta u(r)^{\rm{s,c}}_{\parallel}|^{p}\right\rangle\/,$ (5b)
$\text{and}\quad S_{\rm p}(r)\sim r^{\zeta_{p}}\/,\quad$ $S^{\rm s}_{\rm p}(r)\sim r^{\zeta_{p}^{\rm s}}\/,\quad S^{\rm c}_{\rm p}(r)\sim r^{\zeta_{p}^{\rm c}}\/.$ (5c)
The scaling relations hold for $\eta\ll r\ll L_{\rm I}$, where $\eta$ is the dissipation scale and $L_{\rm I}$ is the integral scale. All the exponents, $\zeta_{p}$, $\zeta_{p}^{\rm s}$ and $\zeta_{p}^{\rm c}$ are nonlinear monotone increasing functions of $p$ (Konstandin et al., 2012; Wang et al., 2017, 2012b) (see Appendix B.2).
3.1 Pair dispersion
Figure 2: Mean-squared separation. (a) Log-log plots of the time evolution of the mean-squared separation, $\left\langle R^{2}(t)\right\rangle$, between two tracers. The growth of $\left\langle R^{2}\right\rangle$ is faster in S1 than in S2, and faster in C2 than in C1. Inset: Semi-log graphs of $\left\langle R^{2}\right\rangle$ for S1 and S2 displaying exponential growth at early times. (b) Log-log plots of $\left\langle R^{2}\right\rangle$ compensated by $t^{2/z}$, where $z=1-\zeta_{\rm 2}/2$; power-law growth of $\left\langle R^{2}\right\rangle$ across its inertial range values appears to occur in every run; in C1, $\left\langle R^{2}\right\rangle$ clearly scales more slowly than $t^{2/z}$ in this range.
In order to study pair dispersion, we investigate how the separation, $R(t)$, between a pair of nearest-neighbor tracers, separated initially by the grid-spacing (approximately $0.3\eta$), evolves with time. We calculate $\left\langle R^{2}(t)\right\rangle$ by averaging, at every instant, over the $2n$ separations. Our results are shown in Fig. (2a). At early times, $\left\langle R^{2}(t)\right\rangle$ grows exponentially for all the runs [For S1 and S2 we show the plots in the inset of Fig. (2a)]. This range of exponential growth is very small for C1 and C2. Similar results (Jullien, 2003) have been obtained for incompressible 2D turbulence. Exponential growth up to a length scale typically implies that the flow across the tracers is smooth up to that scale. This holds for our simulations because at such small time scales most of the pair separations lie away from shocks. We note that at late times the behavior of $\left\langle R^{2}\right\rangle$ is not universal, it depends on precisely how the DNS is forced.
If $R(t)$ lies in the inertial range, then $R^{2}\sim t^{2/(1-h)}$, where $h$ is the local Hölder exponent of the velocity fluctuations in space. If we use Kolmogorov’s simple scaling hypothesis, $\zeta_{\rm 2}=2h$ which yields $\left\langle R^{2}(t)\right\rangle\sim t^{2/z}$, where $z=1-\zeta_{\rm 2}/2$. In incompressible flows, $\zeta_{\rm 2}=2/3$ and $z=2/3$, which corresponds to Richardson scaling. For our simulations, in each case, we calculate $\zeta_{\rm 2}$ from the scaling of the second order structure functions, $S_{\rm 2}(r)$. In Fig. (2b), we display the plots of $\left\langle R^{2}\right\rangle$ compensated by $t^{2/z}$. It appears that there is a range of scales over which $\left\langle R^{2}\right\rangle\sim t^{2/z}$ in S1, S2 and C2, while the increase in $\left\langle R^{2}(t)\right\rangle$ in C1 is visibly slower that $t^{2/z}$. However, in every case, the scaling regime is too small to make robust quantitative measurements of the values of $z$ from the $\left\langle R^{2}(t)\right\rangle$ versus $t$ plots. Hence, we turn to exit-times.
3.2 Exit times
To study the exit-time statistics, we now consider only those tracer pairs whose initial separations, $R$, belong to the inertial range, i.e., $\eta\ll R\ll L_{\rm I}$. Given the periodic boundary conditions, there are $2n$ distinct separations for any $R$. As the tracers evolve with the flow, we calculate the first instances of time at which $R$ decreases below $43R/$ 4 or exceeds $23R/$ 2, denoting them as halving times, $\tau_{\rm H}(R)$, and doubling times, $\tau_{\rm D}(R)$, respectively. We consider 25 different values of $R$, see Appendix C. The statistics of exit times is expected to reveal the dynamic multiscaling exponents if the distances $R$, $23R/$ 2 and $43R/$ 4 lie within the inertial range. As long as this condition is satisfied, we expect our results to be independent of the particular choice of the thresholds, $23R/$ 2 or $43R/$ 4. Crucially, over the inertial range, the equal–time structure functions $S_{\rm p}(r)$, $S^{\rm s}_{\rm p}(r)$ and $S^{\rm c}_{\rm p}(r)$ [see  (5)] show scaling; see Appendix B.2. For each $R$, we calculate the moments $\left\langle\tau_{\rm H}^{-p}(R)\right\rangle$ and $\left\langle\tau_{\rm D}^{-p}(R)\right\rangle$ over three large-eddy-turnover-time. The averaging is over all $2n$ pairs with the same value of $R$. We find that these moments scale with R as,
$\left\langle\tau_{\rm D}^{-p}(R)\right\rangle\sim R^{-\chi^{\rm D}_{\rm p}}\quad\text{and}\quad\left\langle\tau_{\rm H}^{-p}(R)\right\rangle\sim R^{-\chi^{\rm H}_{\rm p}}\,,$ (6)
in the all cases (see Appendix C). Note that the scaling ranges in the different cases are slightly different as they depends on how the flow is stirred. However, in every run, they extend more than a decade in $R$ for all $p$. Moreover, their range of scaling overlaps with with those of the equal-time structure functions (5).
We extract the exponents, $\chi^{\rm D}_{\rm p}$, $\chi^{\rm H}_{\rm p}$, $\zeta_{p}$, $\zeta_{p}^{\rm c}$ and $\zeta_{p}^{\rm s}$, by linear least–square fitting of the log-log plots of $\left\langle\tau_{\rm D}^{-p}(R)\right\rangle$, $\left\langle\tau_{\rm H}^{-p}(R)\right\rangle$, $S_{\rm p}(r)$, $S^{\rm s}_{\rm p}(r)$ and $S^{\rm c}_{\rm p}(r)$, respectively. We repeat the above procedure for an ensemble of $M=10$ snapshots in the statistically stationary state. For each exponent, its final value is its ensemble mean, and its error bars are defined by the standard deviation. In Fig. (3) we plot the dynamic exponents, $\chi^{\rm D}_{\rm p}$ and $\chi^{\rm H}_{\rm p}$, for all the runs, up to $p=6$. Both are nonlinear functions of $p$, indicating dynamic multiscaling. The spreads of the shaded regions in Fig. 3 show the error bars of the dynamic scaling exponents.
Figure 3: Dynamic exponents for doubling and halving times (see text) $\chi^{\rm D}_{\rm p}$ and $\chi^{\rm H}_{\rm p}$, respectively, as functions of the order $p$ which takes integer and fractional values up to $6$. The blue shaded region corresponds to the bridge relation, $p-\zeta_{p}$, derived from the multifractal model. The pink shaded region corresponds to the expression $p-\zeta_{p}^{\rm s}$.
The multifractal model (Parisi & Frisch, 1985; Frisch, 1996) does not distinguish between doubling and halving times, and hence, for both of them it predicts the bridge relation, $\chi^{\rm{H,D}}_{\rm{p}}=p-\zeta_{p}$. Our data for the halving–time exponents $\chi^{\rm H}_{\rm p}$ agree well with this in all the runs (see Fig. 3). However, we find unambiguously that $\chi^{\rm D}_{\rm p}\neq\chi^{\rm H}_{\rm p}$, thereby violating the simple expectation based on the multifractal model.
For the two runs with a solenoidal external force, S1 and S2, we find $\chi^{\rm D}_{\rm p}=p-\zeta_{p}^{\rm s}$, i.e., the bridge relation from the multifractal model holds for the doubling-time exponents, but with $\zeta_{p}$ replaced by $\zeta_{p}^{\rm s}$. Together this demonstrates that for the runs S1 and S2: the solenoidal part of the velocity field is primarily responsible for doubling, but both the solenoidal and irrotational parts affect halving. Note that the bridge relations hold for the two values of Ma we have used although the precise values of the exponents themselves depend on Ma.
For the two irrotationally-forced runs, C1 and C2, besides $\chi^{\rm H}_{\rm p}=p-\zeta_{p}$, we find that the relation $\chi^{\rm H}_{\rm p}=p-\zeta_{p}^{\rm c}$ also holds because the scaling exponents, $\zeta_{p}^{\rm c}$, of the irrotational component of the velocity are quite close to $\zeta_{p}$. Crucially, $\chi^{\rm D}_{\rm p}\neq\chi^{\rm H}_{\rm p}$ again. Furthermore, unlike the runs S1 and S2, $\chi^{\rm D}_{\rm p}\neq p-\zeta_{p}^{\rm s}$ [see Fig. (3)]. This indicates that the growth of $R$ is not determined solely by the solenoidal part of the velocity field, but is affected significantly by the regions of the flows with ${\bm{\nabla}}\cdot\bm{u}>0$. A quantitative understanding of these exponents remains an open question. Finally, given the accuracy of our data, we do not detect any dependence of these exponents on Ma.
Let us now try to identify the structures in the flow that are primarily responsible for the observed scaling by using the multifractal model. We consider the velocity field to comprise an infinity of interwoven sets, $\mathcal{S}_{\rm h}$, each characterized by a scaling exponent, $h$, and fractal dimension, $D(h)$. For any point ${\bm{x}}\in\mathcal{S}_{\rm h}$, $\delta u_{\parallel}(R)\sim R^{h}$, in the limit $R/L_{\rm I}\to 0$. We use this model to calculate the dynamic exponents, $\chi_{\rm{p}}$. The characteristic time scale (here, exit time), $\tau_{\bm{x}}(R)\sim R/[\delta u_{\parallel}(R)]\sim R^{1-h}$. Its moments of negative orders, $-p$ ($p>0$), scale as
[…]
$\implies-\chi_{\rm{p}}=\inf_{h}\left[ph+d-p-D(h)\right]\quad\text{and}\quad D(h)=\inf_{p}\left(ph+d-p+\chi_{\rm{p}}\right)\,.$ (7b)
The spatial dimensionality $d=2$, and $dw(h)$ gives the weights of the different exponents. For $R/L_{\rm I}\to 0$, the integral in  (7a) is evaluated by the method of steepest descents. Equation (7b) implies that $D(h)$ and $\chi_{\rm{p}}$ are Legendre transforms of each other. Physically, $D(h)$ reveals the dimensionality of structures that contribute to the scaling of $\delta u_{\parallel}(R)$ with the exponent $h$, unveiling the multifractal spectrum. Note that, even if $D(h)$ is not concave its Legendre transform $\chi_{\rm{p}}$ will be concave; however, the inverse transform returns not $D(h)$ but its concave hull (see, e.g., Frisch, 1996, section 8.5).
Figure 4: D(h) vs h. Plots of (a) $D_{\rm H}(h)$ and (b) $D_{\rm D}(h)$ versus $h$ for all the runs. $D_{\rm H}(h)$ changes slightly as we change the type of external force and Ma; the dependence of $D_{\rm D}(h)$ on Ma is much stronger in case of the solenoidal external force than the irrotational one. For S2, C1 and C2, $D_{\rm D}(h)$ takes values close to unity but not S1, implying that 1D structures contribute to the scaling of the doubling times in the former cases.
Here, we define $D_{\rm H}(h)$ and $D_{\rm D}(h)$, which are calculated from $\chi^{\rm H}_{\rm p}$ and $\chi^{\rm D}_{\rm p}$, respectively, by using Eq.  (7b). Displayed in Fig. (4), they clearly show that the halving and doubling times are influenced by different structures. In Fig. 4(a), we note that changes in the type of external force and Ma have very little effect on $D_{\rm H}(h)$. In all cases, it goes down to unity as $h\to 0$. Thus, it is reasonable to conclude that the halving times are affected by shocks, which are one-dimensional structures characterized by $h\to 0$. Since $\chi^{\rm H}_{\rm p}=p-\zeta_{p}$, $D_{\rm H}(h)$ overlaps with the multifractal spectrum evaluated from $\zeta_{p}$. The dependence of $D_{\rm D}(h)$ on Ma in S1 and S2 is much stronger, as shown in Fig. (4b). In S1 ($\mbox{Ma}\simeq 1$), $D_{\rm D}(h)$ does not go down to $1$, implying that the structures that influence the increase of $R(t)$ are more space-filling than shocks. In S2 ($\mbox{Ma}\simeq 2.5$), $D_{\rm D}(h)$ does decrease to unity indicating that nearly one-dimensional structures mediate the fast growth of $R$. Since $h\simeq 0.4>0$ when $D_{\rm D}(h)\simeq 1$ and $\chi^{\rm D}_{\rm p}=p-\zeta_{p}^{\rm s}$, we infer that these structures are the thin elongated patches of high vorticity, which appear in the vicinity of the shocks, clearly visible in Fig.1. Such patches do exist for S1, but they are relatively weaker and broader. Finally, we note from Fig. (4b) that the function $D_{\rm D}(h)$ takes values close to unity in the runs, C1 and C2, as well. The development of a quantitative understanding of the exponents, $\chi^{\rm D}_{\rm p}$, in these runs remains a challenge.
4 Conclusion and Discussions
We have thus obtained insights into the multifractal nature of tracer pair dispersion in transonic to supersonic compressible turbulence of an isothermal ideal gas in 2D. The doubling and halving times of pair separations robustly quantify this intermittency and multifractality. Their statistics are different from each other, and the scaling exponents of the moments of their distributions, $\chi^{\rm D}_{\rm p}$ and $\chi^{\rm H}_{\rm p}$, vary nonlinearly with $p$. Therefore, these exit times encapsulate the dynamic multiscaling properties of compressible turbulence. While $\chi^{\rm H}_{\rm p}$ satisfies its multifractal model-based bridge relation and is universal, $\chi^{\rm D}_{\rm p}$ is not, depending on both the external force and the turbulent Mach number, Ma. This non-universality of the inertial-range pair-dispersion statistics is remarkable, given that the inertial-range fluctuations are supposed to be independent of the external force. In the solenoidally-stirred flows, $\chi^{\rm D}_{\rm p}$ is expressible solely in terms of the scaling exponents of the solenoidal velocity fluctuations, implying that predominantly vortical structures influence the fast growth of pair separations. We show that their fractal dimensions vary considerably with Ma which explains the Ma-dependence of $\chi^{\rm D}_{\rm p}$. In contrast, in the irrotationally–stirred flows, pair dispersion is significantly accelerated by the positive flow divergences and $\chi^{\rm D}_{\rm p}$ does not depend on Ma – its comprehensive understanding remains an open problem, whose solution may require that we go beyond the multifractal model.
Note that we have previously undertaken similar studies in simpler models of compressible turbulence – the randomly-forced 1D (De et al., 2023) and 2D (De et al., 2024) Burgers equation. They have a few qualitative similarities with the results here – evidence of dynamic multiscaling and $\chi^{\rm D}_{\rm p}\neq\chi^{\rm H}_{\rm p}$. However, the Burgers equation assumes certain unphysical simplifications: (a) zero pressure (tracers remain trapped on shocks forever), (b) zero vorticity at all times, and (c) bifractal velocity fluctuations – which the model used here does not; thereby displaying richer, more realistic and more complicated features. This additional complexity also precludes the formulation of a heuristic theoretical framework for the exit times which was possible in De et al. (2024).
We expect the key qualitative content of our results to be valid in three dimensions (3D) as well, although 3D, high-Ma compressible turbulence is different from its 2D counterpart. In particular, the robustness and indispensability of the two kinds of exit times in the exploration of the intermittency of pair dispersion, and the nature of their dependence on Ma and the external force, should carry over to 3D. Highly supersonic turbulent flows often govern various astrophysical processes, e.g. star formation in molecular clouds, and the transport and mixing of gases along with their chemical kinetics in the interstellar media (Mac Low & Klessen, 2004; Elmegreen & Scalo, 2004). Our work sets out an effective paradigm which can potentially provide in-depth insights into these aspects.
[…]
[Acknowledgements]DM and RP would like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme Anti-diffusive dynamics: from sub-cellular to astrophysical scales, where partial work on this paper was undertaken. This work was supported by EPSRC grant EP/R014604/1.
[…]
[Funding]DM acknowledges the support of the Swedish Research Council Grant No. 638-2013-9243. Nordita is partially supported by NordForsk. S.D. thanks the PMRF (India) for support. R.P. and S.D. thank the Science and Engineering Research Board (SERB), Anusandhan National Research Foundation (ANRF) and the National Supercomputing Mission (NSM), India for support, and the Supercomputer Education and Research Centre (IISc) for computational resources.
[…]
[Declaration of interests]The authors report no conflict of interest.
[…]
[Data availability statement]The data that support the findings of this study are available from the corresponding author upon request.
[…]
[Author ORCIDs]S. De: https://orcid.org/$0$-0002-1020-968X ; D. Mitra: https://orcid.org/$0$-0003-4861-8152 ; R. Pandit: https://orcid.org/$0$-0003-0948-8834
Appendix A Numerical simulation
In two–dimensions, by taking curl and divergence of the governing equations 2 we obtain two partial differential equations for the two scalars $\omega=\hat{z}\cdot{\bm{\nabla}}\times\bm{u}$ and $\phi={\bm{\nabla}}\cdot\bm{u}$. We solve these two equations by pseudospectral method. We use the second order Adams–Bashforth method for time evolution of all the terms in the PDEs except the external force, which is integrated by the Euler–Marayuma (Higham, 2001) scheme because it is white–in–time.
The velocity field at the off–grid location of the Lagrangian tracers is obtained by using bi-linear interpolation. For the time evolution of the tracers we also use a second order Runge–Kutta scheme.
A.1 Pseudospectral algoithm for compressible flows
The use of pseudospectral methods in incompressible turbulence has been long championed by Orszag and others (Canuto et al., 2007). They have also been used extensively in DNS of compressible turbulence in both two and three dimensions (Kritsuk et al., 2007). A key limitation of the pseudospectral methods for compressible flows is that they are are unable to capture true discontinuties as opposed to, e.g., Reimann solvers. Shocks in pseudospectral solvers have finite width set by the value of the shock viscosity and this width must be resolved. We discuss the role of shock viscosity in greater detail in section A.2. A further drawback is that they are prone to spurious oscillations near the shocks. We filter out such oscillations by appropriate de-aliasing,see section A.3.
In all cases, the Kolmogorov scale, $\eta$, is approximately 66668888 times the grid spacing. Shocks are generally resolved across 444410101010 grid points, which is of the order of $\eta$.
A.2 Shock viscosity
Figure 5: PDF of the shock viscosity, $\zeta_{\rm bulk}$, in the NESS of different runs on a log-log scale. While all values of $\zeta_{\rm bulk}$ are less than $\nu$ in the low-Mach number runs, C1 and S1, a small fraction (about 1%) of them are higher than $\nu$ in the higher-Mach number runs, C2 and S2. These broaden the shocks to a greater extent and enable their resolution at higher Ma.
The shock viscosity, $\zeta_{\rm bulk}$, is a spatially non-uniform bulk viscosity which is locally enhanced in the neighbourhood of shocks and serves to resolve the strong shocks on a finite-resolution grid in the high Mach number simulations. We evaluate $\zeta_{\rm bulk}$ on the basis of the prescription outlined in Ref. Von Neumann & Richtmyer (1950). At each grid $(i,j)\equiv{\bm{x}}$,
$\zeta_{\rm bulk}({\bm{x}})=c_{\rm shock}\left\langle\tilde{\phi}\right\rangle_{5}\delta x\delta y\,,$ (8)
where $\delta x=\delta y=L/N$ are the grid-spacings along each direction, and $\left\langle...\right\rangle_{5}$ denotes the average over a $5\times 5$ domain centred at $(i,j)$.
$\tilde{\phi}=\begin{cases}|{\bm{\nabla}}\cdot\bm{u}|&\text{for}\quad{\bm{\nabla}}\cdot\bm{u}<0\,;\\ 0&\text{otherwise}\,.\end{cases}$ (9)
In S1 and C1, we choose $c_{\rm shock}=1$ while, in S2 and C2, we use $c_{\rm shock}=5$). Thus, $\zeta_{\rm bulk}$ is non-zero around shocks, decreasing quickly with distance from the latter, and is zero elsewhere. In Fig. 5 we show the PDFs of $\zeta_{\rm bulk}$ for the four runs. We do not consider the locations where $\zeta_{\rm bulk}({\bm{x}})=0$, i.e. away from shocks. While $\zeta_{\rm bulk}({\bm{x}})<\nu$ for all ${\bm{x}}$ in the low-Ma runs (C1 and S1), nearly $1\%$ of its values are greater than $\nu$ in the higher-Ma flows, C2 and S2.
A.3 De-aliasing
To remove aliasing errors, we use an exponential filter (Hou & Li, 2007) acting on the nonlinear terms:
$\mathcal{F}(k)=\exp\left[-36.0\left(\frac{k}{k_{\rm m}}\right)^{36}\right]\,,$ (10)
where $k_{\rm m}=N/2$ and $N$ is the number of grid points along either directions. The function, $\mathcal{F}(k)$ is approximately unity up to $k\approx N/3$ and drops quickly but smoothly to almost zero beyond that, hence our aliasing cutoff is approximately $N/3$. We use a smooth filter because it reduces oscillations around the shocks (Hou & Li, 2007).
Appendix B Equal time statistics
Figure 6: (a) The probability distribution functions (PDFs), $\mathcal{P}(\rho)$, of $\rho$ in the different runs; for a given type of external force, the PDF corresponding to the supersonic flow has broader tails than the transonic one. (b) The PDFs, $\mathcal{P}(|c_{\rm s}^{2}{\bm{\nabla}}\ln\rho|)$, of the magnitude of the pressure gradient force per unit mass, $|c_{\rm s}^{2}{\bm{\nabla}}\ln\rho|$, which tends to push tracers away from shocks. The PDFs for the transonic flows have longer tails than the supersonic ones. (c) Complementary cumulative PDFs (CPDFs), $Q_{+}({\bm{\nabla}}\cdot\bm{u})$, of the positive velocity divergences in the irrotationally-driven flows, C1 and C2. The supersonic case, C2, has stronger flow divergences, as evidenced from the broader tail of $Q_{+}$, than the transonic case, C1.
The probability distribution functions (PDFs) of the density, $\rho$, in the NESS of the different runs is shown in Fig. (6a). The PDFs corresponding to the higher-Ma runs (S2 and C2) have broader tails than those of the lower-Ma runs (S1 and C1), indicating a stronger clustering of the tracers in the former. The same may be inferred from Fig. (6b) which displays the PDFs, $\mathcal{P}(|c_{\rm s}^{2}{\bm{\nabla}}\ln\rho|)$, of the pressure gradient force per unit mass that tends to push the tracers out of the clusters. Among the irrotationally-driven flows, C2 has stronger flow divergences away from shocks than C1, as inferred from the complementary cumulative PDFs, $Q_{+}({\bm{\nabla}}\cdot\bm{u})$, of the positive flow divergences in Fig. (6c).
B.1 Energy spectra
Figure 7: Velocity spectra. Log-log plots of the compensated spectra of the total velocity (blue), $E(k)=|\hat{{\bm{u}}}({\bm{k}})|^{2}$, its compressive component (red), $E_{c}(k)=|\hat{\bm{u}^{\rm c}}({\bm{k}})|^{2}$, and its solenoidal components (violet), $E_{s}(k)=|\hat{\bm{u}^{\rm s}}({\bm{k}})|^{2}$, for the different runs; in every image, the dashed line represents $k^{-3}$; every spectrum has been normalized by $\sum_{k}E(k)$ in each of the respective runs
We display the shell-integrated energy spectra for velocity, $E(k)=|\hat{{\bm{u}}}({\bm{k}})|^{2}$, its solenoidal component, $E_{\rm s}(k)=|\hat{\bm{u}^{\rm s}}({\bm{k}})|^{2}$, and its dilational component, $E_{\rm c}(k)=|\hat{\bm{u}^{\rm c}}({\bm{k}})|^{2}$, compensated by $k^{2}$ in Fig. 7. In all of the cases, we observe $E(k)\sim k^{-2}$ and $E_{\rm c}(k)\sim k^{-2}$ for a decade or more. In the irrotationally-forced runs, $E_{\rm c}(k)\gg E_{\rm s}(k)$ across whole of the inertial range – the difference decreases as Ma increases. $E_{\rm s}(k)$, however, does not appear to have a well-defined power-law regime although it does appear to have a small scaling range with $E_{\rm s}(k)\sim k^{-3}$, marked by a dashed line. .
B.2 Equal–time structure functions
We calculate equal-time longitudinal structure functions of the total velocity, $S_{\rm p}(r)$, and its solenoidal, $S^{\rm s}_{\rm p}(r)$, and dilational, $S^{\rm c}_{\rm p}(r)$, see Fig. 8. In each image, the regime of power-law fitting has been demarcated by the dashed lines. The scaling regimes of $S_{\rm p}(r)$, $S^{\rm c}_{\rm p}(r)$ and $S^{\rm s}_{\rm p}(r)$ are not exactly identical – there is partial overlap between them and all of them belong to the regime, $\eta\ll r\ll L_{\rm I}$. The fact that, even at a fixed Reynolds number, the scaling range of different quantities are different is found in incompressbile turbulence (Yakhot & Sreenivasan, 2004; Schumacher et al., 2007) and has also been observed in compressible turbulence Wang et al. (2017). We select the range of scales over which we look for scaling of exit times as the range over which $S_{\rm p}(r)$ shows scaling.
To calculate the scaling exponent $\zeta_{p}$ of $S_{\rm p}(r)$ we proceed the following way. We first consider the data from one single snapshot. Identify a scaling range by eye – this range is typically about a decade, but varies from one scheme of forcing to another. Then fit a straight line to the log-log plot of the structure functions to extract $\zeta_{p}$. We then collect many snapshots into an ensemble. The value of $\zeta_{p}$ we plot in Fig. (9) is the mean over this ensemble and the error is the standard deviation over the same ensemble. In statistical learning theory, this method is called BAGGing or Bootstrap AGGregation (Breiman, 1996; Mehta et al., 2019). The same method is applied to extract $\zeta_{p}^{\rm c}$ and $\zeta_{p}^{\rm s}$. Figure 9 shows that: (a) The solenoidal velocity fluctuations in S1 are less intermittent than those in S2; (b) At large $p$, $\zeta_{p}$ is much closer to $\zeta_{p}^{\rm c}$ because $S_{\rm p}(r)$ is dominated by the contribution from the longitudinal velocity differences across shocks which are predominantly irrotational; and (c) The saturation of $\zeta_{p}$ and $\zeta_{p}^{\rm c}$ at $p\gtrsim 2$ is brought about by the shocks.
Figure 8: Equal-time structure functions. Log-log plots of structure functions: $S^{\rm c}_{\rm p}(r)$ (left column), $S^{\rm s}_{\rm p}(r)$ (middle column) and $S_{\rm p}(r)$ (right column) versus $r$ for $p=2$ (blue), $p=4$ (red) and $p=6$ (violet) in runs (a) S1, (b) S2, (c) C1 and (d) C2. The vertical dashed lines demarcate the region of power-law fitting in each case across which we extract $\zeta_{p}^{\rm c}$, $\zeta_{p}^{\rm s}$ and $\zeta_{p}^{\rm s}$. In each figure, the data has been averaged over a single non-equilibrium steady state (NESS) snapshot.
Figure 9: Equal-time exponents. Scaling exponents, $\zeta_{p}$, $\zeta_{p}^{\rm s}$ and $\zeta_{p}^{\rm c}$ of the structure functions, $S_{\rm p}(r)$, $S^{\rm s}_{\rm p}(r)$ and $S^{\rm c}_{\rm p}(r)$, of different orders, $p$.
Appendix C Lagrangian averages and dynamic exponents
Our data for the moments of exit–times are averaged over $2n$ uniformly distributed pair separations for a single value of the inital separation $R$, with $n=4096^{2}$. Then the calculations are repeated for $25$ values of $R$ within the inertial range. The log-log plot of the moments of exit times as a function of $R$ are shown in Fig. (10). We run the DNS long enough (close to $3$ large–eddy–turnover–time) such that we have more than $10^{7}$ intervals that cross the thresholds for every value of $R$. This process is repeated for $M=10$ times by introducing the tracers to $M$ statistically independent initial conditions. Thus for every $R$ we have at least $10^{8}$ samples. We randomly divide our samples into $M=10$ ensembles and calculate the $M$ individual values of the exponents from each ensemble. The values of the exponents we quote are averaged over the ensemble and the the error we set is the standard deviation of the mean values obtained in each member. In statistical learning theory, this technique is known as BAGGing, or Bootstrap AGGregation (Breiman, 1996; Mehta et al., 2019). Therefore, from the sizes of the errorbars, we can conclude that the Lagrangian averages have converged. To check this further we calculate the exponents from $M/2$ randomly selected members of the ensemble and compare them with the exponents obtained from all ($M$) the members, see Fig. (11). The excellent overlap between the values of the dynamic exponents in these two cases demonstrate the convergence of the Lagrangian averages.
Figure 10: Moments of Lagrangian exit times. Log-log plots of $\left\langle\tau_{\rm D}^{-p}(R)\right\rangle$ and $\left\langle\tau_{\rm H}^{-p}(R)\right\rangle$ versus $R$ for $p=2$ (blue), $p=4$ (red) and $p=6$ (violet) in runs (a) S1, (b) S2, (c) C1 and (d) C2. The vertical dashed lines demarkates the region of power–law fitting.
Figure 11: Convergence of Lagrangian averages. The dynamic exponents calculated using half ($5$) the members of the ensemble, compared with the ones calculated using the whole ensemble ($M=10$), displayed in Fig. 3 of the main text. The excellent overlap between the exponents and their errorbars, in every case affirms the convergence of the Lagrangian averages.

### v3

@@
 Abstract
-Statistical properties of the pair dispersion of Lagrangian particles (tracers) in incompressible turbulent flows provide insights into transport and mixing.
-We explore the same in transonic to supersonic compressible turbulence of an isothermal ideal gas in two dimensions, driven by large-scale solenoidal and irrotational stirring forces, via direct numerical simulations.
-We find that the scaling exponents of the order-$p$ negative moments of the distribution of exit times – in particular, the doubling and halving times of pair separations – are nonlinear functions of $p$.
+We study statistical properties of the pair dispersion of Lagrangian particles in transonic to supersonic compressible turbulence of an isothermal ideal gas in two dimensions, driven by large-scale solenoidal and irrotational stirring forces, via direct numerical simulations.
+We find that the scaling exponents of the order-$p$ negative moments of the doubling and halving times of pair separations are nonlinear functions of $p$.
 Furthermore, the doubling and halving time statistics are different.
 The halving-time exponents are universal – they satisfy their multifractal model-based prediction, irrespective of the nature of the stirring.
-However, the doubling-time exponents are not.
+The doubling-time exponents are not.
 In the solenoidally-stirred flows, the doubling time exponents can be expressed solely in terms of the multifractal scaling exponents obtained from the structure functions of the solenoidal component of the velocity.
-Moreover, they depend strongly on the Mach number, Ma, as elongated patches of high vorticity emerge along shock fronts at high Ma.
+Moreover, they depend strongly on the Mach number, Ma.
 In contrast, in the irrotationally-stirred flows, the doubling-time exponents do not satisfy any known multifractal model-based relation, and are independent of Ma.
-Our findings are of potential relevance to astrophysical disks and molecular clouds wherein turbulent transport and mixing of gases often govern chemical kinetics and the rates of formation of stars and planetesimals.
-keywords:
+Our work is a generalization of Richardson’s law of pair dispersion to compressible turbulence.
 Compressible turbulence; Particle/Fluid Flow
-1 Introduction
-[…]
-We begin by summarizing the evidence in support of Richardson’s law in incompressible turbulence.
-Theoretically, it can be understood as a modeling of the evolution of $R(t)$ by a diffusive process (Richardson, 1926; Benzi & Toschi, 2023) with a diffusion constant, $K(R)\sim R^{4/3}$, that stems (Benzi & Toschi, 2023) from Kolmogorov’s 1941 (K41) theory (Kolmogorov, 1941).
-An alternative interpretation of Richardson’s law is in terms of a dynamic scaling exponent, $z=2/3$, such that typical time scales, $\tau(r)$, of turbulent eddies of size $r$ scale as $\tau(r)\sim r^{z}$.
-It is now well established, from experiments and DNS, that K41 must be generalized to include intermittency, i.e., the $p$-th moment of the Eulerian velocity difference, $\delta u$, across a length scale $r$ (the order-$p$ velocity structure function), scales as $\left\langle\delta u^{p}(r)\right\rangle\sim r^{\zeta_{p}}$, where $\zeta_{p}$ is a nonlinear concave function of $p$ (Frisch, 1996).
-This also dictates that turbulence must, in fact, have an infinite number of dynamic exponents (L’vov et al., 1997; Mitra & Pandit, 2004, 2005; Ray et al., 2008, 2011; Pandit et al., 2008), suggesting the presence of intermittency corrections to Richardson’s law, which itself enjoys feeble support from experiments (Ott & Mann, 2000; Shnapp et al., 2018; Elsinga et al., 2022; Tan & Ni, 2022; Shnapp et al., 2023) and direct numerical simulations (DNS) (Salazar & Collins, 2009; Bitane et al., 2012, 2013; Bragg et al., 2016; Buaria et al., 2015; Elsinga et al., 2022) – the observed scaling range of $\left\langle R^{2}\right\rangle$ is at most a decade.
-A better approach to investigate the intermittency of pair dispersion in turbulence is to use exit times, defined as the time, $\tau_{\alpha}(R)$, it takes for $R$ to cross $\alpha R$ for the first time, where $\alpha$ is fixed.
-The choice $\alpha>1$ corresponds to the doubling times [^1] To be accurate, $\alpha=2$ means “doubling”.
-Nevertheless, in practice, any $\alpha>1$ is used (Boffetta & Sokolov, 2002a), $\tau_{\rm D}$, for which (Boffetta et al., 1999; Boffetta & Sokolov, 2002a, b)
+One of the pillars of our understanding of turbulence is Richardson’s law [1]: $\left\langle R^{2}(t)\right\rangle\sim t^{3}$, where $R$ is the separation between a pair of Lagrangian particles or tracers, $t$ is time, and $\left\langle\cdot\right\rangle$ denotes averaging over the statistically stationary turbulent state.
+This law holds for incompressible flows if $R$ lies within the inertial range of length scales, across which the nonlinear interactions dominate over viscous dissipation.
+Here, we seek a generalization of Richardson’s law to compressible turbulent flows, which are prevalent in many astrophysical systems [2, 3, 4].
+We begin by summarizing the present status of Richardson’s law in incompressible turbulence.
+Theoretically, it can be understood as a modeling of the evolution of $R(t)$ by a diffusive process [1, 5] with a diffusion constant, $K(R)\sim R^{4/3}$, that stems [5] from Kolmogorov’s 1941 (K41) theory [6].
+Stated differently, Richardson’s law implies that there exists a dynamic scaling exponent, $z=2/3$, such that typical time scales, $\tau(r)$, of turbulent eddies of size $r$ scale as $\tau(r)\sim r^{z}$.
+It is now well established, from experiments and direct numerical simulations (DNS), that K41 must be generalized to include intermittency, i.e., the $p$-th moment of the Eulerian velocity difference, $\delta u$, across a length scale $r$ (the order-$p$ velocity structure function), scales as $\left\langle\delta u^{p}(r)\right\rangle\sim r^{\zeta_{p}}$, where $\zeta_{p}$ is a nonlinear concave function of $p$ [7].
+This also dictates that turbulence must, in fact, have an infinite number of dynamic exponents [8, 9, 10, 11, 12, 13], suggesting the presence of intermittency corrections to Richardson’s law, which itself enjoys feeble support from experiments [14, 15, 16, 17, 18] and DNS [19, 20, 21, 22, 23, 16] – the observed scaling range of $\left\langle R^{2}\right\rangle$ is at most a decade.
+A better approach to investigate the intermittency of pair dispersion in turbulence is to use exit times, defined as the time, $\tau_{\sigma}(R)$, it takes for $R$ to cross $\sigma R$ for the first time, where $\sigma$ is fixed.
+The choice $\sigma>1$ corresponds to the doubling times [^1] To be accurate, $\sigma=2$ means “doubling”.
+Nevertheless, in practice, any $\sigma>1$ is used [26], $\tau_{\rm D}$, for which [25, 26, 27]
 $\left\langle\tau_{\rm D}^{-p}(R)\right\rangle\sim R^{-\chi^{\rm D}_{\rm p}}\,,\quad\text{where}\quad\chi^{\rm D}_{\rm p}=p-\zeta_{p}\,.$ (1)
-This bridge relation between the dynamic exponents, $\chi^{\rm D}_{\rm p}$, and $\zeta_{p}$ has been obtained by using the multifractal model of turbulence (Parisi & Frisch, 1985; Frisch, 1996).
-In compressible flows, in addition to doubling times, we need to define halving times, $\tau_{\rm H}(R)$, for which $\alpha<1$ and the corresponding scaling exponents are $\chi^{\rm H}_{\rm p}$.
+This bridge relation between the dynamic exponents, $\chi^{\rm D}_{\rm p}$, and $\zeta_{p}$ has been obtained by using the multifractal model of turbulence.
+In compressible flows, in addition to doubling times, we need to define halving times $\tau_{\rm H}(R)$, for which $\sigma<1$ and the corresponding scaling exponents are $\chi^{\rm H}_{\rm p}$.
 As the multifractal model does not distinguish between the doubling and halving times, $\chi^{\rm H}_{\rm p}$ are also expected naively to satisfy the same bridge relation,  (1).
-Although there have been many studies of the statistics of velocity increments in compressible turbulence (Boldyrev et al., 2002b, a; Schmidt et al., 2008; Konstandin et al., 2012; Wang et al., 2012a; Kritsuk et al., 2007; Wang et al., 2017; Jagannathan & Donzis, 2016), none have investigated the intermittency of tracer pair dispersion.
-We explore this via simulations of forced compressible turbulence of an isothermal ideal gas, governed by the Navier-Stokes equations.
-We restrict ourselves to two spatial dimensions in order to reach high resolutions.
-Two-dimensional (2D) incompressible turbulence is different from its 3D counterpart because, in addition to the kinetic energy, the enstrophy is conserved in the limit of zero viscosity and forcing.
-Theory, experiments and DNS indicate that $\left\langle R^{2}(t)\right\rangle$ grows exponentially with time in the direct–enstrophy–cascade regime (Bennett, 1984; Rivera & Ecke, 2005; Jullien, 2003; LaCasce & Ohlmann, 2003).
-In 2D compressible turbulence, there is small-scale dissipation of kinetic energy via acoustic waves (Falkovich & Kritsuk, 2017).
-It has been shown that $\left\langle R^{2}(t)\right\rangle$ grows as a power–law beyond the dissipation scales in models of such flows, e.g., surface flows (Cressman & Goldburg, 2003; Li et al., 2024) and the 2D Burgers equation (De et al., 2024).
-Here too, we unveil a power-law growth of $\left\langle R^{2}(t)\right\rangle$ across similar spatial scales.
-While, $\chi^{\rm H}_{\rm p}$ obey the multifractal model-based bridge relation and are universal, $\chi^{\rm D}_{\rm p}\neq\chi^{\rm H}_{\rm p}$, in disagreement with the multifractal model.
-If the external force is solenoidal: (a) $\chi^{\rm D}_{\rm p}=p-\zeta_{p}^{\rm s}$, where $\zeta_{p}^{\rm s}$ are the scaling exponents of structure functions of the solenoidal component of the velocity field; and (b) the values of $\chi^{\rm D}_{\rm p}$ (and consequently, $\zeta_{p}^{\rm s}$) depend on Ma.
-This indicates that vortical structures play the dominant role in pair dispersion, and we demonstrate that their fractal dimensions vary significantly with Ma, accounting for the Ma-dependence of $\chi^{\rm D}_{\rm p}$.
+Although there have been many studies of the statistics of velocity increments in compressible turbulence [28, 29, 30, 31, 32, 33] and tracer dispersion in compressible surface flows are known not to obey Richardson’s law [34], the intermittency of tracer pair dispersion have not been studied.
+We explore this via simulations of forced compressible turbulence of an isothermal ideal gas, governed by the Navier-Stokes equations in two dimensions.
+In 2D compressible turbulence forced at an intermediate scale, Refs.
+[35, 36] had obtained two cascades of the kinetic energy: an inverse cascade from the energy injection scale to large length scales and a forward cascade to small length scales.
+In what follows, we choose the energy-injection scales to be close to the linear size of our simulation domain; and we focus on the forward- cascade regime, which we show to be intermittent.
+Figure 1: Typical snapshots of $\nabla\cdot{\bf u}$, the vorticity ${\bm{\omega}}={\bm{\nabla}}\times\bm{u}$, and $\log_{10}\rho$, for solenoidally-driven supersonic turbulence, run S2 (${\bm{\nabla}}\cdot\bm{u}$ and $\bm{\omega}$} are normalized by their respective root-mean-square values).
+Thin filaments in the ${\bm{\nabla}}\cdot\bm{u}$ profile are the shocks.
+The vorticity, $\bm{\omega}$}, is large in the vicinity of curved shocks.
+There are strong density gradients across shocks.
+We unveil a power-law growth of $\left\langle R^{2}(t)\right\rangle$, but the scaling regime is too small to make robust quantitative measurements.
+Thus, we resort to the exit-time approach to calculate the dynamic exponents $\chi^{\rm H}_{\rm p}$ and $\chi^{\rm D}_{\rm p}$.
+We find that $\chi^{\rm H}_{\rm p}$ obey the multifractal model-based bridge relation and are universal, i.e., does not depend on how the turbulence in forced.
+But contrary to the prediction of the multifractal model $\chi^{\rm D}_{\rm p}\neq\chi^{\rm H}_{\rm p}$.
+Moreoever, $\chi^{\rm D}_{\rm p}$ is not universal.
+If the external force is solenoidal, then: (a) $\chi^{\rm D}_{\rm p}=p-\zeta_{p}^{\rm s}$, where $\zeta_{p}^{\rm s}$ are the scaling exponents of structure functions of the solenoidal component of the velocity field; and (b) $\chi^{\rm D}_{\rm p}$ depend on the Mach number (Ma).
 If the force is purely irrotational, $\chi^{\rm D}_{\rm p}$ neither follows any known bridge relation, nor depends on Ma.
-In this case, the regions of large flow divergence affect pair dispersion.
-In Section 2, we describe the model, and outline its parameters and the details of its DNS.
-In Section 3, we delineate our analyses of particle pair-dispersion and Lagrangian exit-times, and then provide the concluding remarks in Section 4.
-2 Model and its numerical simulation
-The equations (Haugen et al., 2022) governing the dynamics of a compressible isothermal ideal gas are:
-$\partial_{t}\rho+\partial_{j}(\rho u_{j})$ $=0\,,$ (2a)
-$\partial_{t}u_{i}+u_{j}\partial_{j}u_{i}$ $=-\frac{1}{\rho}\partial_{i}P+\frac{1}{\rho}\partial_{j}(\nu\rho\sigma_{ij}+\rho\zeta_{\rm bulk}\delta_{ij}\partial_{k}u_{k})+\mathcal{P}_{ij}(\xi)f_{j}\,,$ (2b)
-where $\bm{u}({\bm{x}},t)$ is the velocity field, $\rho({\bm{x}},t)$ is the density field, $P$ is the pressure, $\nu$ is the kinematic shear viscosity, $\zeta_{\rm bulk}$ is the bulk viscosity, $\delta_{ij}$ is the Kronecker delta, $\mathcal{P}_{ij}(\xi)f_{j}$ is the external force, $\sigma_{ij}=(1/2)\left(\partial_{j}u_{i}+\partial_{i}u_{j}\right)-(1/2)\delta_{ij}\partial_{k}u_{k}$ is the traceless rate-of-strain tensor, and the equation of state is given by $P=c_{\rm s}^{2}\rho$, where $c_{\rm s}$ is the (constant) speed of sound.
-The quantity, ${\bm{f}}$, which we call external force or stirring, is white-in-time and its Fourier transform has compact support at $|{\bm{k}}|=k_{\rm f}=3$, where ${\bm{k}}$ is the wave vector.
-The Fourier transform of the projection operator, $\mathcal{P}_{ij}(\xi)$ is, $\hat{\mathcal{P}}_{ij}(\xi)\equiv\xi\delta_{ij}+(1-2\xi)k_{i}k_{j}/k^{2}$.
-We consider two cases, $\xi=0$ and $1$, which correspond, respectively, to an irrotational and a solenoidal external force (see Table 1).
-The assumption of isothermality holds when the medium is optically thin (transparent to radiation) in which case the heat generated by dissipation at shocks is rapidly radiated away.
-This is generally valid for molecular clouds which have low density and solar metallicity (Mac Low & Klessen, 2004).
-We perform pseudospectral DNS of Eq.
-(2) on a 2D doubly-periodic square domain of side $L=2\pi$ with $N^{2}=4096^{2}$ collocation points.
-The bulk viscosity follows the prescription outlined by Von Neumann & Richtmyer (1950): it is zero everywhere except near the shocks.
-We assume $\nu$ to be a constant.
-This is a common practice; see, e.g., Haugen et al.
-(2022); Federrath et al.
-(2011); Richard et al.
+We further identify, through the multifractal model, the dimensions of the structures that play the dominant role in pair dispersion.
+We perform pseudospectral DNS of the equations of compressible, isothermal gas with a random white-in-time external force (at the Fourier modes with $\mid{\bm{k}}\mid=k_{\rm f}=3$) on a 2D doubly-periodic square domain of side $L=2\pi$ with $N^{2}=4096^{2}$ collocation points.
+The assumption of isothermality holds when the gas is optically thin (transparent to radiation) in which case the heat generated by the dissipation at shocks is rapidly radiated away.
+This is generally valid for molecular clouds which have low density and solar metallicity [2].
+In addition to shear viscosity, we use a bulk viscosity which is zero everywhere except near the shocks [37].
+Shocks are generally resolved across $4$ to $10$ grid points.
+After the flow reaches a statistically stationary state we seed it uniformly with $n=N^{2}$ Lagrangian particles and track them as the flow evolves.
+They cluster at the shocks and their number density is proportional to $\rho({\bm{x}},t)$.
+Using the Helmholtz decomposition, we write $\bm{u}=\bm{u}^{\rm s}+\bm{u}^{\rm c}$, where $\bm{u}^{\rm s}$ and $\bm{u}^{\rm c}$ are, respectively, the solenoidal ($\nabla\cdot\bm{u}^{\rm s}=0$) and irrotational ($\nabla\times\bm{u}^{\rm c}=0$) components of the flow.
+The relative importance of $\bm{u}^{\rm s}$ and $\bm{u}^{\rm c}$ in the total kinetic energy is measured by the solenoidal ratio, $\Psi\equiv\left\langle\rho|\bm{u}^{\rm s}|^{2}\right\rangle/\left\langle\rho|{\bm{u}}|^{2}\right\rangle$.
+See Appendix A for further details on the numerical method.
+We perform four sets of DNSs, S1 and S2 for solenoidal external force and C1 and C2 for irrotational external force, see Table 1 for all the parameters of the runs.
+Typically, subsonic and supersonic compressible flow can be qualitatively different.
+Hence we perform a set of simulations that are transsonic, Runs S1 and C1, $\mbox{Ma}\simeq 1$, and another set S2 and C2 are supersonic, $\mbox{Ma}>1$.
+In Fig.
+(1) we show representative pseudocolour plots of ${\bm{\nabla}}\cdot\bm{u}$, ${\bm{\omega}}={\bm{\nabla}}\times\bm{u}$, and $\log_{10}\rho$ for the run S2; see Fig.
+(3) for representative snapshots The shocks are clearly visible as filamentary structures with large negative values of ${\bm{\nabla}}\cdot\bm{u}$.
+Naively, we may expect that the presence of the shocks would not show up in pseudocolour plot of $\omega$, but this not the case because although a shock front that is a straight line does not contribute to vorticity, curved shock fronts do.
+In our simulations, (a) the density gradients are larger for the supersonic cases compared to the transonic ones, and (b) the vorticity is more uniformly distributed in the transonic cases than the supersonic ones.
+We can define the $p$-th order longitudinal equal-time structures functions of $\bm{u}$ as well as its components, $\bm{u}^{\rm s}$ and $\bm{u}^{\rm c}$:
+$\delta u_{\parallel}(r)\equiv\left[\bm{u}({\bm{x}}+\bm{r})-\bm{u}({\bm{x}})\right]\cdot\frac{\bm{r}}{r}\/,$ (2a)
+$\text{and}\quad\delta u_{\parallel}^{\rm{s,c}}(r)\equiv\left[\bm{u}^{\rm{s,c}}({\bm{x}}+\bm{r})-\bm{u}^{\rm{s,c}}({\bm{x}})\right]\cdot\frac{\bm{r}}{r}\/;$ (2b)
+$\text{with}\quad S_{\rm p}(r)=\left\langle|\delta u(r)_{\parallel}|^{p}\right\rangle\/,\quad S_{\rm p}^{\rm s,c}(r)=\left\langle|\delta u(r)^{\rm{s,c}}_{\parallel}|^{p}\right\rangle\/,$ (2c)
+$\text{and}\quad S_{\rm p}(r)\sim r^{\zeta_{p}}\/,\quad S^{\rm s}_{\rm p}(r)\sim r^{\zeta_{p}^{\rm s}}\/,\quad S^{\rm c}_{\rm p}(r)\sim r^{\zeta_{p}^{\rm c}}\/.$ (2d)
+The scaling relations hold for $\eta\ll r\ll L_{\rm I}$, where $\eta$ is the energy dissipation scale.
+All the equal-time exponents, $\zeta_{p}$, $\zeta_{p}^{\rm s}$ and $\zeta_{p}^{\rm c}$ are nonlinear monotone increasing functions of $p$ [29, 32, 31].
+The spectra of turbulent fluctuations and the equal-time structures functions are shown in Appendices B.2 and B.3 respectively.
+We calculate $\left\langle R^{2}(t)\right\rangle$ by averaging, at every instant, over the $2n$ separations where $n=N^{2}$ is the number of tracers.
+See Fig.
+(8) for plots of $\left\langle R^{2}(t)\right\rangle$ versus time for all the runs.
+At early times, $\left\langle R^{2}(t)\right\rangle$ grows exponentially for all the runs.
+This range of exponential growth is very small for C1 and C2.
+Similar results [38] have been obtained for incompressible 2D turbulence.
+Exponential growth up to a length scale typically implies that the flow across is smooth up to that scale.
+This holds for our simulations because at such small time scales most of the pair separations lie away from shocks.
+At late times the behaviour of $\left\langle R^{2}\right\rangle$ is not universal, it depends on precisely how the DNS is forced.
+If $R(t)$ lies in the inertial range, then $R^{2}\sim t^{2/(1-h)}$, where $h$ is the local Hölder exponent of the velocity fluctuations in space.
+If we use Kolmogorov’s simple scaling hypothesis, $\zeta_{\rm 2}=2h$ which yields $\left\langle R^{2}(t)\right\rangle\sim t^{2z}$, where $z=1/(1-\zeta_{\rm 2}/2)$.
+In incompressible flows $\zeta_{\rm 2}=2/3$ and $z=3/2$ that corresponds to Richardson’s law.
+We calculate $\zeta_{\rm 2}$ from the scaling of the second order structure functions, $S_{\rm 2}(r)$.
+We find that there is a range of scales over which $\left\langle R^{2}\right\rangle\sim t^{2z}$ in S1, S2 and C2; although not in C1.
+Crucially, in every case, the scaling regime is too small to make robust quantitative measurements of the values of $z$ from the $\left\langle R^{2}(t)\right\rangle$ versus $t$ plots.
+Hence, we turn to exit-times.
+Figure 2: Dynamic exponents for doubling and halving times (see text) $\chi^{\rm D}_{\rm p}$ and $\chi^{\rm H}_{\rm p}$, respectively, as functions of the order $p$ for the two supersonic runs [S2 (left panel) and C2 (right panel)].
+The blue shaded region corresponds to the bridge relation, $p-\zeta_{p}$, derived from the multifractal model.
+The pink shaded region corresponds to the expression $p-\zeta_{p}^{\rm s}$.
+In both S2 and C2, the multifractal prediction for $\chi^{\rm H}_{\rm p}$ is correct within errorbars, but not for $\chi^{\rm D}_{\rm p}$.
+For S2 $\chi^{\rm D}_{\rm p}=p-\zeta_{p}^{\rm s}$ is correct within errorbars.
+For C2 $\chi^{\rm D}_{\rm p}$ does not satisfy any known bridge relation.
+In the stationary state, we choose $2n$ tracer pairs whose relative distance $R$ lies in the inertial range and then calculate $\tau_{\rm H}(R)$ and $\tau_{\rm D}(R)$ as the first instant at which $R$ goes below $3R/4$ or above $3R/2$, respectively  [^2] To calculate $\left\langle R^{2}(t)\right\rangle$ we use initial values of $R$ to be approximately one grid spacing.
+But to calculate exit time we must start with $R$ within the inertial range.
+The statistics of exit–times is expected to show scaling (or multiscaling) if the distances $R$, $3R/2$ and $3R/4$ lie within the inertial range  [^3] As long as this condition is satisfied, we expect our results to be independent of the particular choice of the thresholds, $3R/2$ or $3R/4$..
+Specifically, the inertial range is defined as the range of scales over which the equal–time structure functions $S_{\rm p}(r)$, $S^{\rm s}_{\rm p}(r)$ and $S^{\rm c}_{\rm p}(r)$ [see  (2)] show scaling.
+We find that the negative moments of the exit-time distributions scale as
+$\left\langle\tau_{\rm D}^{-p}(R)\right\rangle\sim R^{-\chi^{\rm D}_{\rm p}}\quad\text{and}\quad\left\langle\tau_{\rm H}^{-p}(R)\right\rangle\sim R^{-\chi^{\rm H}_{\rm p}}\,.$ (3)
+over almost a decade in $R$.
+We extract the exponents, $\chi^{\rm D}_{\rm p}$, $\chi^{\rm H}_{\rm p}$, $\zeta_{p}$, $\zeta_{p}^{\rm c}$ and $\zeta_{p}^{\rm s}$, by linear least–square fitting of the log-log plots of $\left\langle\tau_{\rm D}^{-p}(R)\right\rangle$, $\left\langle\tau_{\rm H}^{-p}(R)\right\rangle$, $S_{\rm p}(r)$, $S^{\rm s}_{\rm p}(r)$ and $S^{\rm c}_{\rm p}(r)$, respectively.
+In Fig.
+(2) we plot the dynamic exponents, $\chi^{\rm D}_{\rm p}$ and $\chi^{\rm H}_{\rm p}$, for the runs S2 and C2, up to $p=6$.
+Both are nonlinear functions of $p$, indicating dynamic multiscaling.
+The spreads of the shaded regions in Fig.
+2 show the error bars of the $\zeta_{p}$’s and $\zeta_{p}^{\rm s}$’s.
+The moments of exits times and the dynamic exponents calculated from them for all the runs are shown in Appendix B.5.
+See Appendix B.7 for our method of estimating error.
+The multifractal model [41, 7] does not distinguish between doubling and halving times, and hence, for both of them it predicts the same bridge relation, $\chi^{\rm{H,D}}_{\rm{p}}=p-\zeta_{p}$.
+This bridge relation agrees well with our data, for the halving times, (see Fig.
+2), but not for the doubling times.
+We find unambiguously that $\chi^{\rm D}_{\rm p}\neq\chi^{\rm H}_{\rm p}$, thereby violating the simple expectation based on the multifractal model.
+For the two runs with a solenoidal external force, S1 and S2, we find $\chi^{\rm D}_{\rm p}=p-\zeta_{p}^{\rm s}$, i.e., the bridge relation from the multifractal model holds for the doubling-time exponents, but with $\zeta_{p}$ replaced by $\zeta_{p}^{\rm s}$.
+Together this demonstrates that, for the runs S1 and S2, (a) the solenoidal part of the velocity field is primarily responsible for the increase of $R$, and (b) both the solenoidal and irrotational parts affect the decrease of $R$.
+Note that the bridge relations hold for the two values of Ma we have used although the precise values of the exponents themselves depend on Ma.
+For the two irrotationally-forced runs, C1 and C2, besides $\chi^{\rm H}_{\rm p}=p-\zeta_{p}$, we find that the relation $\chi^{\rm H}_{\rm p}=p-\zeta_{p}^{\rm c}$ also holds because the scaling exponents, $\zeta_{p}^{\rm c}$, of the irrotational component of the velocity are quite close to $\zeta_{p}$.
+Crucially, $\chi^{\rm D}_{\rm p}\neq\chi^{\rm H}_{\rm p}$ again.
+Furthermore, unlike the runs S1 and S2, $\chi^{\rm D}_{\rm p}\neq p-\zeta_{p}^{\rm s}$ (see Fig.
+(2) and Fig.
+(10)).
+This indicates that the growth of $R$ is not determined solely by the solenoidal part of the velocity field, but is affected significantly by the regions of the flows with ${\bm{\nabla}}\cdot\bm{u}>0$.
+A quantitative understanding of these exponents remains an open question.
+Finally, given the accuracy of our data, we do not detect any dependence of these exponents on Ma.
+In the multifractal model, the velocity field is constructed from an infinity of interwoven sets each characterised by a scaling exponent $h$ and corresponding fractal dimension $D(h)$.
+It is straightforward to show (see Appendix B.6 ) that the dynamic exponents $\chi_{\rm{p}}$ and $D(h)$ are Legendre transforms of each other:
+$\chi_{\rm{p}}=-\inf_{h}\left[ph+d-p-D(h)\right]\/$ (4a)
+$\text{and}\quad D(h)=\inf_{p}\left(ph+d-p+\chi_{\rm{p}}\right)\/.$ (4b)
+Thus from $\chi_{\rm{p}}$ we determine $D(h)$ – the dimensionality of structures that contribute to dynamic scaling [^4] Even if $D(h)$ is not concave its Legendre transform $\chi_{\rm{p}}$ is concave; however, the inverse transform returns not $D(h)$ but its concave hull [see, e.g., 7, section 8.5]..
+In addition, we must define $D_{\rm H}(h)$ and $D_{\rm D}(h)$, which are calculated from $\chi^{\rm H}_{\rm p}$ and $\chi^{\rm D}_{\rm p}$, respectively, by using Eq.
+(4b).
+This reveals that the halving and doubling times are influenced by structures of different dimensionality.
+All the multifractal spectra are shown in Fig.
 (11).
-After the flow reaches a statistically stationary state we seed it uniformly with $n=N^{2}$ Lagrangian particles or tracers at every grid point, such that the distance between a pair of nearest-neighbor tracers is the grid spacing.
+We find that external force and Ma have very little effect on $D_{\rm H}(h)$ – in all our runs, $D_{\rm H}(h)\to 1$ as $h\to 0$.
+Thus the halving times are controlled by shocks, which are one-dimensional structures characterized by $h\to 0$.
+As expected, $D_{\rm H}(h)$ overlaps with the multifractal spectrum evaluated from equal time exponents $\zeta_{p}$.
+In contrast, the function $D_{\rm D}(h)$ is not universal.
+In S1 ($\mbox{Ma}\simeq 1$), $D_{\rm D}(h)$ does not go down to $1$, implying that the structures that influence the scaling of $R(t)$ are more space-filling than shocks.
+In S2 ($\mbox{Ma}\simeq 2.5$), we find $D_{\rm D}(h)\simeq 1$ for $h\simeq 0.4>0$ we infer that these structures are not shocks but thin elongated patches of high vorticity, which appear in the vicinity of the shocks, clearly visible in Fig.1.
+Such patches do exist for S1, but they are relatively weaker and broader.
+Finally, we note that the function $D_{\rm D}(h)$ takes values close to unity in the runs, C1 and C2, as well.
+Note that we have previously undertaken similar studies in simpler models of compressible turbulence – the randomly-forced Burgers equation in one [43] and two dimensions [44].
+They have a few qualitative similarities with the results here – evidence of dynamic multiscaling and $\chi^{\rm D}_{\rm p}\neq\chi^{\rm H}_{\rm p}$.
+There are few key differences between the Burgers equation and the present, more realistic model: (a) zero pressure (tracers remain trapped on shocks forever), (b) zero vorticity at all times, and (c) bifractal velocity fluctuations.
+These additional complexity also precludes the formulation of building a theoretical framework for the exit times which was possible in De et al.
+[44].
+In summary, in trying to generalize Richardson’s law of pair dispersion to compressible flows, we learn three important lessons: (a) Compressible turbulence in two dimensions show dynamic multiscaling; (b) The dynamic multiscaling obtained here is qualitatively different from its incompressible counterpart; we need to consider both doubling and halving times and the two have different dynamic multiscaling exponents.
+The exponents $\chi^{\rm H}_{\rm p}$ satisfies a multifractal model–based bridge relation and is universal but $\chi^{\rm D}_{\rm p}$ is not.
+They depend on both the external force and the turbulent Mach number, Ma.
+This non-universality of the inertial–range pair–dispersion statistics is remarkable, given that the inertial-range fluctuations are supposed to be independent of the external force.
+(c) This dynamic multiscaling cannot be understood within the framework of the present multifractal model of turbulence.
+Consequently, a clear understanding of statistics of pair dispersion in compressible turbulence remains an open problem, whose solution may require that we go beyond the multifractal model.
+Furthermore, note that in incompressible turbulence, the Richardson dispersion can be understood as a scale dependent effective viscosity and this serves as a measure of mixing by turbulent flows.
+We find that in compressible turbulence this description is fundamentally flawed because we have both doubling and halving times which have different scaling properties.
+As turbulence in most astrophysical context is compressible, mixing in large scale astrophysical systems, e.g., molecular clouds, must be revisited in the light of our results.
+We speculate that the key qualitative content of our results would be valid in three dimensions (3D) as well.
+In particular, the necessity of the two kinds of exit times, dynamic multiscaling, and intermittency of pair dispersion should carry over to 3D.
+Highly supersonic turbulent flows govern various astrophysical processes, e.g.
+star formation in molecular clouds, and the transport and mixing of gases along with their chemical kinetics in the interstellar media [2, 4].
+Our work sets out an effective paradigm which can potentially provide in-depth insights into these aspects.
+Acknowledgements.
+DM and RP thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme Anti-diffusive dynamics: from sub-cellular to astrophysical scales, where partial work on this paper was undertaken.
+This work was supported by EPSRC grant EP/R014604/1.
+DM acknowledges the support of the Swedish Research Council Grant No.
+638-2013-9243.
+Nordita is partially supported by NordForsk.
+S.D.
+thanks the PMRF (India) for support.
+R.P.
+and S.D.
+thank the Science and Engineering Research Board (SERB), Anusandhan National Research Foundation (ANRF) and the National Supercomputing Mission (NSM), India for support, and the Supercomputer Education and Research Centre (IISc) for computational resources.
+Appendix A Numerical simulation
+The governing equations, in vectorial notation, are as follows:
+$\frac{\partial\rho}{\partial t}+{\bm{\nabla}}\cdot(\rho\bm{u})=0\,$, (5a)
+$\frac{\partial\bm{u}}{\partial t}+\bm{u}\cdot{\bm{\nabla}}\bm{u}=[2\nu{\bm{S}}+{\bm{I}}_{2}(\zeta_{\rm bulk}{\bm{\nabla}}\cdot\bm{u}-c_{\rm s}^{2})]\cdot{\bm{\nabla}}(\ln\rho)+\nu{\bm{\nabla}}^{2}\bm{u}+{\bm{\nabla}}\cdot(\zeta_{\rm bulk}{\bm{\nabla}}\cdot\bm{u})+{\bm{f}}_{\rm ext}\,.$ (5b)
+Here $\bm{u}({\bm{x}},t)$ is the velocity field, $\rho({\bm{x}},t)$ is the density field, $P$ is the pressure, $\nu$ is the kinematic shear viscosity taken to be a constant [45, 46, 47].
+$\zeta_{\rm bulk}$ is the bulk viscosity, $\delta_{ij}$ is the Kronecker delta, $S_{ij}\equiv(1/2)\left(\partial_{j}u_{i}+\partial_{i}u_{j}\right)-(1/2)\delta_{ij}\partial_{k}u_{k}$ is the traceless rate-of-strain tensor, and the equation of state is given by $P=c_{\rm s}^{2}\rho$, where $c_{\rm s}$ is the (constant) speed of sound.
+Here, $f_{{\rm ext}\,,i}=\mathcal{P}_{ij}f_{j}$ and ${\bm{I}}_{2}$ is the $2\times 2$ identity matrix.
+In 2D, Eq.
+(5b) can be rewritten in terms of two scalars – the vorticity, $\omega=\hat{z}\cdot{\bm{\nabla}}\times\bm{u}$, and the divergence of velocity, $\mathcal{D}={\bm{\nabla}}\cdot\bm{u}$.
+We use a pseudo-spectral code to solve  (5b) rewritten in $\omega$ and $\mathcal{D}$.
+A.1 De-aliasing
+To remove aliasing errors, we use an exponential filter [48]:
+$\mathcal{F}(k)=\exp\left[-36.0\left(\frac{k}{k_{\rm m}}\right)^{36.0}\right]\,$, (6)
+where $k_{\rm m}=N/2$ is the highest allowed wave-number in the simulation, $N$ the number of grid points in a given dimension.
+The filter $\mathcal{F}(k)$ is approximately unity up to $k\approx N/3$ and drops sharply but smoothly to almost zero beyond that, essentially all modes with $k>N/3$ are strongly damped.
+It is a better choice than the Galerkin-based discrete cut-off filter which are usually used for incompressible turbulence because it reduces oscillations around shocks which result from Galerkin-trucation, enabling us to simulate supersonic turbulence at high Mach numbers while maintaining relatively large Reynolds numbers.
+A.2 External force
+The quantity, ${\bm{f}}({\bm{x}},t)$, is a white-in-time random noise. Its spectrum in the spatial domain is restricted to large spatial scales (small $k$). In our simulations, we use
+$\widehat{{\bm{f}}}_{k_{x}}({\bm{k}},t_{n})$ $=A\cos(\theta_{{\bm{k}},n})e^{i\phi_{{\bm{k}},n}}\/,\$(7a)
+$\text{and}\quad\widehat{{\bm{f}}}_{k_{y}}({\bm{k}},t_{n})$ $=A\sin(\theta_{{\bm{k}},n})e^{i\phi_{{\bm{k}},n}}\/,$ (7b)
+where the amplitude,
+$A=\begin{cases}f_{0}\sqrt{\delta t}&\text{for}\quad k=k_{\rm f}\,;\\ 0&\text{otherwise}\,.\end{cases}$ (8)
+$f_{0}$ is constant, and $\theta_{{\bm{k}},n}$ and $\phi_{{\bm{k}},n}$ are uniformly distributed random numbers belonging to the interval $0,2\pi]$.
+We choose $k_{\rm f}=3$.
+The ratio of power being injected into the solenoidal and irrotational modes is tuned by applying the following projection operator,
+$\hat{\mathcal{P}}_{ij}=\xi\delta_{ij}+(1-2\xi)k_{i}k_{j}/k^{2}\,$, (9)
+which yields ${\bm{f}}_{\rm ext}({\bm{x}},t)$.
+A.3 The shock viscosity $\zeta_{\rm bulk}$
+The shock viscosity, $\zeta_{\rm bulk}$, is a spatially non-uniform bulk viscosity which is locally enhanced in the neighbourhood of shocks and serves to resolve the strong shocks on a finite-resolution grid in the high Mach number simulations.
+We evaluate $\zeta_{\rm bulk}$ on the basis of the prescription outlined in Ref.
+[37].
+At each grid $(i,j)\equiv{\bm{x}}$,
+$\zeta_{\rm bulk}({\bm{x}})=c_{\rm shock}\left\langle\overline{\mathcal{D}}\right\rangle_{5}\delta x\delta y\,$, (10)
+where $\delta x=\delta y=L/N$ are the grid-spacings along each direction, and $\left\langle...\right\rangle_{5}$ denotes the average over a $5\times 5$ domain centred at $(i,j)$.
+$\overline{\mathcal{D}}=\begin{cases}|{\bm{\nabla}}\cdot\bm{u}|&\text{for}\quad{\bm{\nabla}}\cdot\bm{u}<0\,;\\ 0&\text{otherwise}\,.\end{cases}$ (11)
+We choose $c_{\rm shock}\sim\mathcal{O}(1)$. Thus, $\zeta_{\rm bulk}$ is non-zero around shocks, decreasing quickly with distance from them, and zero elsewhere.
+A.4 Timestepping
+We implement the white-in-time external force with Euler Marayuma scheme [49] and exponential Adams-Bashforth method [50] for timestepping the rest of the equation.
+A.5 Lagrangian particle tracking
 The position of the $\rm{m}$-th tracer, $\bm{X}^{\rm m}(t)$, evolves according to
-$\frac{d\bm{X}^{\rm m}(t)}{dt}={\bm{V}}^{\rm m}(t)\quad\text{where}\quad{\bm{V}}^{\rm m}(t)=\bm{u}({\bm{x}},t)\delta({\bm{x}}-\bm{X}^{\rm m})\/.$ (3)
-They cluster at the shocks and their number density maps $\rho({\bm{x}},t)$.
-The Reynolds and turbulent Mach numbers are defined, respectively, as $\mbox{Re}\equiv u_{\rm rms}L/\nu$ and $\mbox{Ma}\equiv u_{\rm rms}/c_{\rm s}$, where $u_{\rm rms}$ is the root-mean-square velocity.
-Using Helmholtz decomposition, we write $\bm{u}=\bm{u}^{\rm s}+\bm{u}^{\rm c}$, where $\bm{u}^{\rm s}$ and $\bm{u}^{\rm c}$ are, respectively, the solenoidal (${\bm{\nabla}}\cdot\bm{u}^{\rm s}=0$) and irrotational (×${\bm{\nabla}}\times\bm{u}^{\rm c}=0$) components of the flow.
-The relative importance of $\bm{u}^{\rm s}$ and $\bm{u}^{\rm c}$ in the total kinetic energy is measured by the solenoidal ratio,
-$\Psi\equiv\frac{\left\langle\rho|\bm{u}^{\rm s}|^{2}\right\rangle}{\left\langle\rho|{\bm{u}}|^{2}\right\rangle}\/,$ (4)
-which is the ratio of the total kinetic energy in the solenoidal modes to the total kinetic energy.
-All these parameters of our DNS are summarized in Table 1.
-We use $L$ as our unit of length, and $T_{\rm E}$, the large-eddy turnover time, as our unit of time.
-Further details about the DNS are given in Appendix A.
+$\frac{d\bm{X}^{\rm m}(t)}{dt}$ $={\bm{V}}^{\rm m}(t)$ (12a)
+$\text{where}\quad{\bm{V}}^{\rm m}(t)$ $=\bm{u}({\bm{x}},t)\delta({\bm{x}}-\bm{X}^{\rm m})\/.$ (12b)
+The velocity at the off-grid position of a tracer is obtained by linear interpolation from the neighboring grid points and the timestepping is done by a Runge-Kutta scheme.
+Appendix B Results
+B.1 Parameters and Snapshots
+The key parameters of our DNS are summarized in Table 1.
+In Fig.
+3 we show representative snapshots of ${\bm{\nabla}}\cdot\bm{u}$, ${\bm{\omega}}={\bm{\nabla}}\times\bm{u}$, and $\log_{10}\rho$ for all runs.
+The shocks are clearly visible as filamentary structures with large negative values of ${\bm{\nabla}}\cdot\bm{u}$.
+Naively, we may expect that the presence of the shocks would not show up in pseudocolour plot of $\bm{\omega}$}, but this not the case – the shocks can be seen as almost one-dimensional manifolds with large absolute values of the vorticity: This is because, although a shock front that is a straight line does not contribute to vorticity, curved shock fronts do.
+In our simulations, (a) the density fluctuations are larger for the supersonic cases compared to the transonic ones and (b) the vorticity is more uniformly distributed in the transonic cases than the supersonic ones, for a given type of external forcing, as observed in Fig.
+3.
 Run $\xi$ $c_{\rm s}$ $u_{\rm rms}$ Ma Re $L_{\rm I}$ $T_{\rm E}$ $\Psi$ $\epsilon$ $\lambda$ $\mbox{Re}_{\lambda}$ $\eta$
-S1 1 0.05 0.053 $1.05$ $6.7\times 10^{3}$ 1.66 32 $0.82$ $8.1\times 10^{-6}$ 0.29 306 0.0111
-S2 1 0.02 0.050 $2.5$ $6.3\times 10^{3}$ 1.20 24 $0.68$ $9.6\times 10^{-6}$ 0.26 255 0.0107
-C1 0 0.05 0.038 $0.75$ $4.7\times 10^{3}$ 1.04 28 $0.19$ $1.3\times 10^{-5}$ 0.16 123 0.0099
-C2 0 0.02 0.040 $2.0$ $5.0\times 10^{3}$ 1.17 29 $0.34$ $8.7\times 10^{-6}$ 0.21 172 0.0109
+S1 1 0.05 0.053 $1.05$ $6.7\times 10^{3}$ 1.66 32 $0.82$ $8.1\times 10^{-6}$ 0.29 306 0.01
+S2 1 0.02 0.050 $2.5$ $6.3\times 10^{3}$ 1.20 24 $0.68$ $9.6\times 10^{-6}$ 0.26 255 0.01
+C1 0 0.05 0.038 $0.75$ $4.7\times 10^{3}$ 1.04 28 $0.19$ $1.3\times 10^{-5}$ 0.16 123 0.01
+C2 0 0.02 0.040 $2.0$ $5.0\times 10^{3}$ 1.17 29 $0.34$ $8.7\times 10^{-6}$ 0.21 172 0.01
 Table 1: Simulation parameters.
@@
 Note that the length scales in this Table are not normalized by $L=2\pi$.
-3 Results
-Figure 1: Typical snapshots of $\nabla\cdot{\bf u}$, vorticity ${\bm{\omega}}=\hat{z}\cdot{\bm{\nabla}}\times\bm{u}$, and $\log_{10}\rho$, in the cases (a) S1, (b) S2, (c) C1 and (d) C2.
-In each plot, ${\bm{\nabla}}\cdot\bm{u}$ and ${\bm{\omega}}$ are normalized by their respective root-mean-square (rms) values.
+Figure 3: Typical snapshots of $\nabla\cdot{\bf u}$, vorticity $\omega=\hat{z}\cdot{\bm{\nabla}}\times\bm{u}$, and $\log_{10}\rho$, in the cases (a) S1, (b) S2, (c) C1 and (d) C2.
+In each plot, ${\bm{\nabla}}\cdot\bm{u}$ and $\omega$ are normalized by their respective root-mean-square (rms) values.
 Shocks are clearly visible in the ${\bm{\nabla}}\cdot\bm{u}$ profiles as the blue filament–like structures.
-The vorticity field, $\omega$, is smoother in S1 than in S2 which contains nearly one-dimensional structures of intense vorticity bordering the shocks; in all figures, the areas of high vorticity lie in the vicinity of shocks.
-Similarly the spatial distribution of $\omega$ is more homogeneous in C1 than in C2.
+The vorticity field, $\omega$, is smoother in S1 than in S2 which contains nearly one-dimensional structures of intense vorticity bordering the shocks; in all figures, the areas of high vorticity lie in the vicinity of shocks Similarly the spatial distribution of $\omega$ is more homogeneous in C1 than in C2.
 The values of $\rho$ are clearly enhanced near the shocks; for a given type of external forcing, $\rho$ tends to take much higher values at higher Mach numbers, Ma.
-We perform four sets of DNSs, two for $\xi=1$, S1 and S2, and two for $\xi=0$, C1 and C2.
-Typically, subsonic and supersonic compressible flow can be qualitatively different.
-Hence we perform a set of simulations that are transonic, Runs S1 and C1, $\mbox{Ma}\simeq 1$, and another set S2 and C2 that are supersonic, $\mbox{Ma}>1$.
-In Fig.
-1 we show pseudocolour plots of ${\bm{\nabla}}\cdot\bm{u}$, ${\bm{\omega}}={\bm{\nabla}}\times\bm{u}$, and $\log_{10}\rho$ for all the cases or runs.
-The shocks are clearly visible as filamentary structures with large negative values of ${\bm{\nabla}}\cdot\bm{u}$.
-Naively, we may expect that the presence of the shocks would not show up in pseudocolour plot of ${\bm{\omega}}$, but this not the case – the shocks can be seen as almost one-dimensional manifolds with large absolute values of the vorticity: This is because, although a shock front that is a straight line does not contribute to vorticity, curved shock fronts do.
-In our simulations, (a) the density fluctuations are larger for the supersonic cases compared to the transonic ones and (b) the vorticity is more uniformly distributed in the transonic cases than the supersonic ones, for a given type of external forcing, as observed in Fig.
-1.
+B.2 cascade and spectra
+Figure 4: Kinetic energy flux for the four DNS runs. We obtain a range of direct cascade.
+Figure 5: Compensated velocity spectra Log-log plots of the compensated (by $k^{2}$) spectra of the total velocity (blue), $E(k)$, its compressive component (red), $E_{\rm c}(k)$, and its solenoidal components (violet), $E_{\rm s}(k)$ for the different runs.
+The spectra are normalized by $\sum_{k}E(k)$.
+Using the Helmholtz decomposition, we write $\bm{u}=\bm{u}^{\rm s}+\bm{u}^{\rm c}$, where $\bm{u}^{\rm s}$ and $\bm{u}^{\rm c}$ are, respectively, the solenoidal ($\nabla\cdot\bm{u}^{\rm s}=0$) and irrotational ($\nabla\times\bm{u}^{\rm c}=0$) components of the flow.
+The relative importance of $\bm{u}^{\rm s}$ and $\bm{u}^{\rm c}$ in the total kinetic energy is measured by the solenoidal ratio,
+$\Psi\equiv\frac{\left\langle\rho|\bm{u}^{\rm s}|^{2}\right\rangle}{\left\langle\rho|{\bm{u}}|^{2}\right\rangle}\/,$ (13)
+which is the ratio of the total kinetic energy in the solenoidal modes to the total kinetic energy.
+In two-dimensional incompressible turbulence, forcing at an intermediate scale gives rise to two cascades: (a) an inverse cascade to large scales and (b) a direct cascade to small scales.
+The statistics in the inverse cascade is close to Gaussian, whereas the statistics in the direct cascade is intermittent.
+In two dimensional compressible turbulence, forced at intermediate scales, kinetic energy shows a dual cascade [35].
+We choose to force at the large scales, $k_{\rm f}=3$, such that we obtain a large range of direct cascade, see Fig.
+(4).
+The statistics of velocity fluctuations in this range is intermittent, as we show later by evaluating the scaling exponents of the $p$-th order structure functions.
+We display the three shell-averaged velocity spectra,
+$E(k)$ $=|\hat{{\bm{u}}}({\bm{k}})|^{2}\/,$ (14a)
+$E_{\rm s}(k)$ $=|\hat{\bm{u}^{\rm s}}({\bm{k}})|^{2}\/,$ (14b)
+$\text{and}\quad E_{\rm c}(k)$ $=|\hat{\bm{u}^{\rm c}}({\bm{k}})|^{2}$ (14c)
+respective for the total, solenoidal, and compressive velocity.
+They are plotted in Fig.
+(5).
+In all the runs, $E_{\rm c}\sim k^{-2}$ over a significant range of scales.
+Earlier works [35] have also observed this over a smaller range of scales and have identified it with acoustic turbulence.
+In S1, $E_{\rm s}(k)$ falls off much faster than that in S2.
+Thus, the small-scale fluctuations of $\bm{u}^{\rm s}$ are much smoother in S1 than those in S2, and this is a reflection of the elongated intense vorticity patches, which are aplenty in S2, but not in S1.
+In S2, $E_{\rm s}(k)$ appears to scale as $k^{-3}$, across a small regime of $k>k_{\rm f}$.
+In the irrotationally-forced runs, $E_{\rm c}(k)\gg E_{s}(k)$ across whole of the inertial range with the difference being small at higher Ma.
+$E_{\rm s}(k)$, however, does not appear to have a well-defined power-law regime.
+B.3 Equal-time structure functions
+Figure 6: Equal-time structure functions.
+Log-log plots of $S_{\rm p}(r)$, $S^{\rm s}_{\rm p}(r)$ and $S^{\rm c}_{\rm p}(r)$, defined in  (15) versus $r$ for $p=2$ (blue), $p=4$ (red) and $p=6$ (violet) in runs (a) S1, (b) S2, (c) C1 and (d) C2.
+The vertical dashed lines demarcate the region of power-law fitting in each case across which we extract $\zeta_{p}$, $\zeta_{p}^{\rm s}$ and $\zeta_{p}^{\rm c}$.
+In each figure, the data has been averaged over a single non-equilibrium steady state (NESS) snapshot.
+Figure 7: Equal-time exponents $\zeta_{p}$ for the four runs.
 In compressible turbulence, we can define $p$-th order longitudinal structures functions of $\bm{u}$ as well as its components, $\bm{u}^{\rm s}$ and $\bm{u}^{\rm c}$:
@@
 The scaling relations hold for $\eta\ll r\ll L_{\rm I}$, where $\eta$ is the dissipation scale and $L_{\rm I}$ is the integral scale.
-All the exponents, $\zeta_{p}$, $\zeta_{p}^{\rm s}$ and $\zeta_{p}^{\rm c}$ are nonlinear monotone increasing functions of $p$ (Konstandin et al., 2012; Wang et al., 2017, 2012b) (see Appendix B.2).
-3.1 Pair dispersion
+All the exponents, $\zeta_{p}$, $\zeta_{p}^{\rm s}$ and $\zeta_{p}^{\rm c}$ are nonlinear monotone increasing functions of $p$ [29, 31, 32].
+We show plots of various equal-time longitudinal structure functions of the total velocity, $S_{\rm p}(r)$, and its solenoidal, $S^{\rm s}_{\rm p}(r)$, and irrotational, $S^{\rm c}_{\rm p}(r)$, components in Fig.
+(6).
+In each image, the regime of power-law fitting has been demarcated by dashed lines.
+In Figs.
+6(a) and 6(b), note that the scaling regime of $S^{\rm s}_{\rm p}(r)$ in S2 occurs across higher values of $r$ than that in S1.
+Moreover, for a given $p$, the slope in the scaling-range of $S^{\rm s}_{\rm p}(r)$ in the former is visibly smaller than that of $S^{\rm s}_{\rm p}(r)$ in the latter.
+For $S_{\rm 2}^{\rm s}(r)$, the nature of this change can also be inferred from the scaling of $E_{\rm s}(k)$ in Fig.
+(5a) and Fig.
+(5()b).
+This is indicative of the appearance of the strip-like intense vorticity patches in S2.
+B.4 Pair dispersion
 Figure 8: Mean-squared separation.
@@
 (b) Log-log plots of $\left\langle R^{2}\right\rangle$ compensated by $t^{2/z}$, where $z=1-\zeta_{\rm 2}/2$; power-law growth of $\left\langle R^{2}\right\rangle$ across its inertial range values appears to occur in every run; in C1, $\left\langle R^{2}\right\rangle$ clearly scales more slowly than $t^{2/z}$ in this range.
-In order to study pair dispersion, we investigate how the separation, $R(t)$, between a pair of nearest-neighbor tracers, separated initially by the grid-spacing (approximately $0.3\eta$), evolves with time.
-We calculate $\left\langle R^{2}(t)\right\rangle$ by averaging, at every instant, over the $2n$ separations.
+We calculate how the separation, $R(t)$, between a pair of nearest-neighbor tracers, separated initially by the grid-spacing (approximately $0.3\eta$), evolves with time.
+We obtain $\left\langle R^{2}(t)\right\rangle$ by averaging, at every instant, over the $2n$ separations.
 Our results are shown in Fig.
@@
 This range of exponential growth is very small for C1 and C2.
-Similar results (Jullien, 2003) have been obtained for incompressible 2D turbulence.
+Similar results [38] have been obtained for incompressible 2D turbulence.
 Exponential growth up to a length scale typically implies that the flow across the tracers is smooth up to that scale.
 This holds for our simulations because at such small time scales most of the pair separations lie away from shocks.
-We note that at late times the behavior of $\left\langle R^{2}\right\rangle$ is not universal, it depends on precisely how the DNS is forced.
+We note that at late times the behavior of $\left\langle R^{2}\right\rangle$ is not universal.
 If $R(t)$ lies in the inertial range, then $R^{2}\sim t^{2/(1-h)}$, where $h$ is the local Hölder exponent of the velocity fluctuations in space.
@@
 However, in every case, the scaling regime is too small to make robust quantitative measurements of the values of $z$ from the $\left\langle R^{2}(t)\right\rangle$ versus $t$ plots.
-Hence, we turn to exit-times.
-3.2 Exit times
-To study the exit-time statistics, we now consider only those tracer pairs whose initial separations, $R$, belong to the inertial range, i.e., $\eta\ll R\ll L_{\rm I}$.
-Given the periodic boundary conditions, there are $2n$ distinct separations for any $R$.
-As the tracers evolve with the flow, we calculate the first instances of time at which $R$ decreases below $43R/$ 4 or exceeds $23R/$ 2, denoting them as halving times, $\tau_{\rm H}(R)$, and doubling times, $\tau_{\rm D}(R)$, respectively.
-We consider 25 different values of $R$, see Appendix C.
-The statistics of exit times is expected to reveal the dynamic multiscaling exponents if the distances $R$, $23R/$ 2 and $43R/$ 4 lie within the inertial range.
-As long as this condition is satisfied, we expect our results to be independent of the particular choice of the thresholds, $23R/$ 2 or $43R/$ 4.
-Crucially, over the inertial range, the equal–time structure functions $S_{\rm p}(r)$, $S^{\rm s}_{\rm p}(r)$ and $S^{\rm c}_{\rm p}(r)$ [see  (5)] show scaling; see Appendix B.2.
-For each $R$, we calculate the moments $\left\langle\tau_{\rm H}^{-p}(R)\right\rangle$ and $\left\langle\tau_{\rm D}^{-p}(R)\right\rangle$ over three large-eddy-turnover-time.
-The averaging is over all $2n$ pairs with the same value of $R$.
-We find that these moments scale with R as,
-$\left\langle\tau_{\rm D}^{-p}(R)\right\rangle\sim R^{-\chi^{\rm D}_{\rm p}}\quad\text{and}\quad\left\langle\tau_{\rm H}^{-p}(R)\right\rangle\sim R^{-\chi^{\rm H}_{\rm p}}\,,$ (6)
-in the all cases (see Appendix C).
-Note that the scaling ranges in the different cases are slightly different as they depends on how the flow is stirred.
-However, in every run, they extend more than a decade in $R$ for all $p$.
-Moreover, their range of scaling overlaps with with those of the equal-time structure functions (5).
-We extract the exponents, $\chi^{\rm D}_{\rm p}$, $\chi^{\rm H}_{\rm p}$, $\zeta_{p}$, $\zeta_{p}^{\rm c}$ and $\zeta_{p}^{\rm s}$, by linear least–square fitting of the log-log plots of $\left\langle\tau_{\rm D}^{-p}(R)\right\rangle$, $\left\langle\tau_{\rm H}^{-p}(R)\right\rangle$, $S_{\rm p}(r)$, $S^{\rm s}_{\rm p}(r)$ and $S^{\rm c}_{\rm p}(r)$, respectively.
-We repeat the above procedure for an ensemble of $M=10$ snapshots in the statistically stationary state.
-For each exponent, its final value is its ensemble mean, and its error bars are defined by the standard deviation.
-In Fig.
-(3) we plot the dynamic exponents, $\chi^{\rm D}_{\rm p}$ and $\chi^{\rm H}_{\rm p}$, for all the runs, up to $p=6$.
-Both are nonlinear functions of $p$, indicating dynamic multiscaling.
-The spreads of the shaded regions in Fig.
-3 show the error bars of the dynamic scaling exponents.
-Figure 3: Dynamic exponents for doubling and halving times (see text) $\chi^{\rm D}_{\rm p}$ and $\chi^{\rm H}_{\rm p}$, respectively, as functions of the order $p$ which takes integer and fractional values up to $6$.
+B.5 Exit times
+The graphs showing the scaling of the moments of the Lagrangian exit times, $\left\langle\tau_{\rm D}^{-p}(r)\right\rangle$ and $\left\langle\tau_{\rm H}^{-p}(r)\right\rangle$, are given in Fig. 9.
+Figure 9: Moments of Lagrangian exit times.
+Log-log plots of $\left\langle\tau_{\rm D}^{-p}(r)\right\rangle$ and $\left\langle\tau_{\rm H}^{-p}(r)\right\rangle$ versus $r$ for $p=2$ (blue), $p=4$ (red) and $p=6$ (violet) in runs (a) S1, (b) S2, (c) C1 and (d) C2.
+The vertical dashed lines demarcate the region of power-law fitting in each case across which we extract $\chi^{\rm H}_{\rm p}$ and $\chi^{\rm D}_{\rm p}$.
+Figure 10: Dynamic exponents for doubling and halving times $\chi^{\rm D}_{\rm p}$ and $\chi^{\rm H}_{\rm p}$, respectively, as functions of the order $p$ which takes integer and fractional values up to $6$.
 The blue shaded region corresponds to the bridge relation, $p-\zeta_{p}$, derived from the multifractal model.
 The pink shaded region corresponds to the expression $p-\zeta_{p}^{\rm s}$.
-The multifractal model (Parisi & Frisch, 1985; Frisch, 1996) does not distinguish between doubling and halving times, and hence, for both of them it predicts the bridge relation, $\chi^{\rm{H,D}}_{\rm{p}}=p-\zeta_{p}$.
-Our data for the halving–time exponents $\chi^{\rm H}_{\rm p}$ agree well with this in all the runs (see Fig.
-3).
-However, we find unambiguously that $\chi^{\rm D}_{\rm p}\neq\chi^{\rm H}_{\rm p}$, thereby violating the simple expectation based on the multifractal model.
-For the two runs with a solenoidal external force, S1 and S2, we find $\chi^{\rm D}_{\rm p}=p-\zeta_{p}^{\rm s}$, i.e., the bridge relation from the multifractal model holds for the doubling-time exponents, but with $\zeta_{p}$ replaced by $\zeta_{p}^{\rm s}$.
-Together this demonstrates that for the runs S1 and S2: the solenoidal part of the velocity field is primarily responsible for doubling, but both the solenoidal and irrotational parts affect halving.
-Note that the bridge relations hold for the two values of Ma we have used although the precise values of the exponents themselves depend on Ma.
-For the two irrotationally-forced runs, C1 and C2, besides $\chi^{\rm H}_{\rm p}=p-\zeta_{p}$, we find that the relation $\chi^{\rm H}_{\rm p}=p-\zeta_{p}^{\rm c}$ also holds because the scaling exponents, $\zeta_{p}^{\rm c}$, of the irrotational component of the velocity are quite close to $\zeta_{p}$.
-Crucially, $\chi^{\rm D}_{\rm p}\neq\chi^{\rm H}_{\rm p}$ again.
-Furthermore, unlike the runs S1 and S2, $\chi^{\rm D}_{\rm p}\neq p-\zeta_{p}^{\rm s}$ [see Fig.
-(3)].
-This indicates that the growth of $R$ is not determined solely by the solenoidal part of the velocity field, but is affected significantly by the regions of the flows with ${\bm{\nabla}}\cdot\bm{u}>0$.
-A quantitative understanding of these exponents remains an open question.
-Finally, given the accuracy of our data, we do not detect any dependence of these exponents on Ma.
-Let us now try to identify the structures in the flow that are primarily responsible for the observed scaling by using the multifractal model.
-We consider the velocity field to comprise an infinity of interwoven sets, $\mathcal{S}_{\rm h}$, each characterized by a scaling exponent, $h$, and fractal dimension, $D(h)$.
-For any point ${\bm{x}}\in\mathcal{S}_{\rm h}$, $\delta u_{\parallel}(R)\sim R^{h}$, in the limit $R/L_{\rm I}\to 0$.
-We use this model to calculate the dynamic exponents, $\chi_{\rm{p}}$.
-The characteristic time scale (here, exit time), $\tau_{\bm{x}}(R)\sim R/[\delta u_{\parallel}(R)]\sim R^{1-h}$.
-Its moments of negative orders, $-p$ ($p>0$), scale as
-[…]
-$\implies-\chi_{\rm{p}}=\inf_{h}\left[ph+d-p-D(h)\right]\quad\text{and}\quad D(h)=\inf_{p}\left(ph+d-p+\chi_{\rm{p}}\right)\,.$ (7b)
-The spatial dimensionality $d=2$, and $dw(h)$ gives the weights of the different exponents.
-For $R/L_{\rm I}\to 0$, the integral in  (7a) is evaluated by the method of steepest descents.
-Equation (7b) implies that $D(h)$ and $\chi_{\rm{p}}$ are Legendre transforms of each other.
-Physically, $D(h)$ reveals the dimensionality of structures that contribute to the scaling of $\delta u_{\parallel}(R)$ with the exponent $h$, unveiling the multifractal spectrum.
-Note that, even if $D(h)$ is not concave its Legendre transform $\chi_{\rm{p}}$ will be concave; however, the inverse transform returns not $D(h)$ but its concave hull (see, e.g., Frisch, 1996, section 8.5).
-Figure 4: D(h) vs h.
+B.6 The multifractal spectrum
+Figure 11: Multifractal spectrum.
 Plots of (a) $D_{\rm H}(h)$ and (b) $D_{\rm D}(h)$ versus $h$ for all the runs.
@@
 For S2, C1 and C2, $D_{\rm D}(h)$ takes values close to unity but not S1, implying that 1D structures contribute to the scaling of the doubling times in the former cases.
-Here, we define $D_{\rm H}(h)$ and $D_{\rm D}(h)$, which are calculated from $\chi^{\rm H}_{\rm p}$ and $\chi^{\rm D}_{\rm p}$, respectively, by using Eq.
-(7b).
-Displayed in Fig.
-(4), they clearly show that the halving and doubling times are influenced by different structures.
-In Fig.
-4(a), we note that changes in the type of external force and Ma have very little effect on $D_{\rm H}(h)$.
-In all cases, it goes down to unity as $h\to 0$.
-Thus, it is reasonable to conclude that the halving times are affected by shocks, which are one-dimensional structures characterized by $h\to 0$.
-Since $\chi^{\rm H}_{\rm p}=p-\zeta_{p}$, $D_{\rm H}(h)$ overlaps with the multifractal spectrum evaluated from $\zeta_{p}$.
-The dependence of $D_{\rm D}(h)$ on Ma in S1 and S2 is much stronger, as shown in Fig.
-(4b).
-In S1 ($\mbox{Ma}\simeq 1$), $D_{\rm D}(h)$ does not go down to $1$, implying that the structures that influence the increase of $R(t)$ are more space-filling than shocks.
-In S2 ($\mbox{Ma}\simeq 2.5$), $D_{\rm D}(h)$ does decrease to unity indicating that nearly one-dimensional structures mediate the fast growth of $R$.
-Since $h\simeq 0.4>0$ when $D_{\rm D}(h)\simeq 1$ and $\chi^{\rm D}_{\rm p}=p-\zeta_{p}^{\rm s}$, we infer that these structures are the thin elongated patches of high vorticity, which appear in the vicinity of the shocks, clearly visible in Fig.1.
-Such patches do exist for S1, but they are relatively weaker and broader.
-Finally, we note from Fig.
-(4b) that the function $D_{\rm D}(h)$ takes values close to unity in the runs, C1 and C2, as well.
-The development of a quantitative understanding of the exponents, $\chi^{\rm D}_{\rm p}$, in these runs remains a challenge.
-4 Conclusion and Discussions
-We have thus obtained insights into the multifractal nature of tracer pair dispersion in transonic to supersonic compressible turbulence of an isothermal ideal gas in 2D.
-The doubling and halving times of pair separations robustly quantify this intermittency and multifractality.
-Their statistics are different from each other, and the scaling exponents of the moments of their distributions, $\chi^{\rm D}_{\rm p}$ and $\chi^{\rm H}_{\rm p}$, vary nonlinearly with $p$.
-Therefore, these exit times encapsulate the dynamic multiscaling properties of compressible turbulence.
-While $\chi^{\rm H}_{\rm p}$ satisfies its multifractal model-based bridge relation and is universal, $\chi^{\rm D}_{\rm p}$ is not, depending on both the external force and the turbulent Mach number, Ma.
-This non-universality of the inertial-range pair-dispersion statistics is remarkable, given that the inertial-range fluctuations are supposed to be independent of the external force.
-In the solenoidally-stirred flows, $\chi^{\rm D}_{\rm p}$ is expressible solely in terms of the scaling exponents of the solenoidal velocity fluctuations, implying that predominantly vortical structures influence the fast growth of pair separations.
-We show that their fractal dimensions vary considerably with Ma which explains the Ma-dependence of $\chi^{\rm D}_{\rm p}$.
-In contrast, in the irrotationally–stirred flows, pair dispersion is significantly accelerated by the positive flow divergences and $\chi^{\rm D}_{\rm p}$ does not depend on Ma – its comprehensive understanding remains an open problem, whose solution may require that we go beyond the multifractal model.
-Note that we have previously undertaken similar studies in simpler models of compressible turbulence – the randomly-forced 1D (De et al., 2023) and 2D (De et al., 2024) Burgers equation.
-They have a few qualitative similarities with the results here – evidence of dynamic multiscaling and $\chi^{\rm D}_{\rm p}\neq\chi^{\rm H}_{\rm p}$.
-However, the Burgers equation assumes certain unphysical simplifications: (a) zero pressure (tracers remain trapped on shocks forever), (b) zero vorticity at all times, and (c) bifractal velocity fluctuations – which the model used here does not; thereby displaying richer, more realistic and more complicated features.
-This additional complexity also precludes the formulation of a heuristic theoretical framework for the exit times which was possible in De et al.
+Following the multifractal model we consider the velocity field to comprise an infinity of interwoven sets, $\mathcal{S}_{\rm h}$, each characterized by a scaling exponent, $h$, and fractal dimension, $D(h)$.
+For any point ${\bm{x}}\in\mathcal{S}_{\rm h}$, $\delta u_{\parallel}(R)\sim R^{h}$, in the limit $R/L_{\rm I}\to 0$.
+We use this model to calculate the dynamic exponents, $\chi_{\rm{p}}$.
+The characteristic time scale (here, exit-time), $\tau_{\bm{x}}(R)\sim R/[\delta u_{\parallel}(R)]\sim R^{1-h}$.
+Its moments of negative orders, p-p ($p>0$), scale as
+$\left\langle\tau^{-p}(R)\right\rangle\sim\int dw(h)\left(\frac{R}{L_{\rm I}}\right)^{d-D(h)-p(1-h)}\sim R^{-\chi_{\rm{p}}}\/;$ (16a)
+$\implies-\chi_{\rm{p}}=\inf_{h}\left[ph+d-p-D(h)\right]\/$ (16b)
+$\text{and}\quad D(h)=\inf_{p}\left(ph+d-p+\chi_{\rm{p}}\right)\/.$ (16c)
+The spatial dimensionality $d=2$, and $dw(h)$ gives the weights of the different exponents.
+For $R/L_{\rm I}\to 0$, the integral in  (16a) is evaluated by the method of steepest descents.
+Equation (16c) implies that $D(h)$ and $\chi_{\rm{p}}$ are Legendre transforms of each other.
+Physically, $D(h)$ reveals the dimensionality of structures that contribute to the scaling of $\delta u_{\parallel}(R)$ with the exponent $h$, unveiling the multifractal spectrum.
+Note that, even if $D(h)$ is not concave its Legendre transform $\chi_{\rm{p}}$ will be concave; however, the inverse transform returns not $D(h)$ but its concave hull [see, e.g., 7, section 8.5].
+B.7 Error analysis
+For equal-time structure functions we take $M=10$, well-separated-in-time, snapshots, from the stationary state of our DNS.
+From each snapshot we calculate the structure function  (15), choose a scaling range by eye estimation (this range is the same is all the snapshots) and fit the log-log plot of the data with a linear function to obtain the equal time scaling exponents.
+In the worst case we obtain close to a decade of clean scaling.
+The structure functions from one such snapshot is shown in Fig.
 (6).
-We expect the key qualitative content of our results to be valid in three dimensions (3D) as well, although 3D, high-Ma compressible turbulence is different from its 2D counterpart.
-In particular, the robustness and indispensability of the two kinds of exit times in the exploration of the intermittency of pair dispersion, and the nature of their dependence on Ma and the external force, should carry over to 3D.
-Highly supersonic turbulent flows often govern various astrophysical processes, e.g.
-star formation in molecular clouds, and the transport and mixing of gases along with their chemical kinetics in the interstellar media (Mac Low & Klessen, 2004; Elmegreen & Scalo, 2004).
-Our work sets out an effective paradigm which can potentially provide in-depth insights into these aspects.
-[…]
-[Acknowledgements]DM and RP would like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme Anti-diffusive dynamics: from sub-cellular to astrophysical scales, where partial work on this paper was undertaken.
-This work was supported by EPSRC grant EP/R014604/1.
-[…]
-[Funding]DM acknowledges the support of the Swedish Research Council Grant No.
-638-2013-9243.
-Nordita is partially supported by NordForsk.
-S.D.
-thanks the PMRF (India) for support.
-R.P.
-and S.D.
-thank the Science and Engineering Research Board (SERB), Anusandhan National Research Foundation (ANRF) and the National Supercomputing Mission (NSM), India for support, and the Supercomputer Education and Research Centre (IISc) for computational resources.
-[…]
-[Declaration of interests]The authors report no conflict of interest.
-[…]
-[Data availability statement]The data that support the findings of this study are available from the corresponding author upon request.
-[…]
-[Author ORCIDs]S. De: https://orcid.org/$0$-0002-1020-968X ; D. Mitra: https://orcid.org/$0$-0003-4861-8152 ; R. Pandit: https://orcid.org/$0$-0003-0948-8834
-Appendix A Numerical simulation
-In two–dimensions, by taking curl and divergence of the governing equations 2 we obtain two partial differential equations for the two scalars $\omega=\hat{z}\cdot{\bm{\nabla}}\times\bm{u}$ and $\phi={\bm{\nabla}}\cdot\bm{u}$.
-We solve these two equations by pseudospectral method.
-We use the second order Adams–Bashforth method for time evolution of all the terms in the PDEs except the external force, which is integrated by the Euler–Marayuma (Higham, 2001) scheme because it is white–in–time.
-The velocity field at the off–grid location of the Lagrangian tracers is obtained by using bi-linear interpolation. For the time evolution of the tracers we also use a second order Runge–Kutta scheme.
-A.1 Pseudospectral algoithm for compressible flows
-The use of pseudospectral methods in incompressible turbulence has been long championed by Orszag and others (Canuto et al., 2007).
-They have also been used extensively in DNS of compressible turbulence in both two and three dimensions (Kritsuk et al., 2007).
-A key limitation of the pseudospectral methods for compressible flows is that they are are unable to capture true discontinuties as opposed to, e.g., Reimann solvers.
-Shocks in pseudospectral solvers have finite width set by the value of the shock viscosity and this width must be resolved.
-We discuss the role of shock viscosity in greater detail in section A.2.
-A further drawback is that they are prone to spurious oscillations near the shocks.
-We filter out such oscillations by appropriate de-aliasing,see section A.3.
-In all cases, the Kolmogorov scale, $\eta$, is approximately 66668888 times the grid spacing. Shocks are generally resolved across 444410101010 grid points, which is of the order of $\eta$.
-A.2 Shock viscosity
-Figure 5: PDF of the shock viscosity, $\zeta_{\rm bulk}$, in the NESS of different runs on a log-log scale.
-While all values of $\zeta_{\rm bulk}$ are less than $\nu$ in the low-Mach number runs, C1 and S1, a small fraction (about 1%) of them are higher than $\nu$ in the higher-Mach number runs, C2 and S2.
-These broaden the shocks to a greater extent and enable their resolution at higher Ma.
-The shock viscosity, $\zeta_{\rm bulk}$, is a spatially non-uniform bulk viscosity which is locally enhanced in the neighbourhood of shocks and serves to resolve the strong shocks on a finite-resolution grid in the high Mach number simulations.
-We evaluate $\zeta_{\rm bulk}$ on the basis of the prescription outlined in Ref.
-Von Neumann & Richtmyer (1950).
-At each grid $(i,j)\equiv{\bm{x}}$,
-$\zeta_{\rm bulk}({\bm{x}})=c_{\rm shock}\left\langle\tilde{\phi}\right\rangle_{5}\delta x\delta y\,,$ (8)
-where $\delta x=\delta y=L/N$ are the grid-spacings along each direction, and $\left\langle...\right\rangle_{5}$ denotes the average over a $5\times 5$ domain centred at $(i,j)$.
-$\tilde{\phi}=\begin{cases}|{\bm{\nabla}}\cdot\bm{u}|&\text{for}\quad{\bm{\nabla}}\cdot\bm{u}<0\,;\\ 0&\text{otherwise}\,.\end{cases}$ (9)
-In S1 and C1, we choose $c_{\rm shock}=1$ while, in S2 and C2, we use $c_{\rm shock}=5$).
-Thus, $\zeta_{\rm bulk}$ is non-zero around shocks, decreasing quickly with distance from the latter, and is zero elsewhere.
-In Fig.
-5 we show the PDFs of $\zeta_{\rm bulk}$ for the four runs.
-We do not consider the locations where $\zeta_{\rm bulk}({\bm{x}})=0$, i.e.
-away from shocks.
-While $\zeta_{\rm bulk}({\bm{x}})<\nu$ for all ${\bm{x}}$ in the low-Ma runs (C1 and S1), nearly $1\%$ of its values are greater than $\nu$ in the higher-Ma flows, C2 and S2.
-A.3 De-aliasing
-To remove aliasing errors, we use an exponential filter (Hou & Li, 2007) acting on the nonlinear terms:
-$\mathcal{F}(k)=\exp\left[-36.0\left(\frac{k}{k_{\rm m}}\right)^{36}\right]\,,$ (10)
-where $k_{\rm m}=N/2$ and $N$ is the number of grid points along either directions.
-The function, $\mathcal{F}(k)$ is approximately unity up to $k\approx N/3$ and drops quickly but smoothly to almost zero beyond that, hence our aliasing cutoff is approximately $N/3$.
-We use a smooth filter because it reduces oscillations around the shocks (Hou & Li, 2007).
-Appendix B Equal time statistics
-Figure 6: (a) The probability distribution functions (PDFs), $\mathcal{P}(\rho)$, of $\rho$ in the different runs; for a given type of external force, the PDF corresponding to the supersonic flow has broader tails than the transonic one.
-(b) The PDFs, $\mathcal{P}(|c_{\rm s}^{2}{\bm{\nabla}}\ln\rho|)$, of the magnitude of the pressure gradient force per unit mass, $|c_{\rm s}^{2}{\bm{\nabla}}\ln\rho|$, which tends to push tracers away from shocks.
-The PDFs for the transonic flows have longer tails than the supersonic ones.
-(c) Complementary cumulative PDFs (CPDFs), $Q_{+}({\bm{\nabla}}\cdot\bm{u})$, of the positive velocity divergences in the irrotationally-driven flows, C1 and C2.
-The supersonic case, C2, has stronger flow divergences, as evidenced from the broader tail of $Q_{+}$, than the transonic case, C1.
-The probability distribution functions (PDFs) of the density, $\rho$, in the NESS of the different runs is shown in Fig.
-(6a).
-The PDFs corresponding to the higher-Ma runs (S2 and C2) have broader tails than those of the lower-Ma runs (S1 and C1), indicating a stronger clustering of the tracers in the former.
-The same may be inferred from Fig.
-(6b) which displays the PDFs, $\mathcal{P}(|c_{\rm s}^{2}{\bm{\nabla}}\ln\rho|)$, of the pressure gradient force per unit mass that tends to push the tracers out of the clusters.
-Among the irrotationally-driven flows, C2 has stronger flow divergences away from shocks than C1, as inferred from the complementary cumulative PDFs, $Q_{+}({\bm{\nabla}}\cdot\bm{u})$, of the positive flow divergences in Fig.
-(6c).
-B.1 Energy spectra
-Figure 7: Velocity spectra.
-Log-log plots of the compensated spectra of the total velocity (blue), $E(k)=|\hat{{\bm{u}}}({\bm{k}})|^{2}$, its compressive component (red), $E_{c}(k)=|\hat{\bm{u}^{\rm c}}({\bm{k}})|^{2}$, and its solenoidal components (violet), $E_{s}(k)=|\hat{\bm{u}^{\rm s}}({\bm{k}})|^{2}$, for the different runs; in every image, the dashed line represents $k^{-3}$; every spectrum has been normalized by $\sum_{k}E(k)$ in each of the respective runs
-We display the shell-integrated energy spectra for velocity, $E(k)=|\hat{{\bm{u}}}({\bm{k}})|^{2}$, its solenoidal component, $E_{\rm s}(k)=|\hat{\bm{u}^{\rm s}}({\bm{k}})|^{2}$, and its dilational component, $E_{\rm c}(k)=|\hat{\bm{u}^{\rm c}}({\bm{k}})|^{2}$, compensated by $k^{2}$ in Fig.
-7.
-In all of the cases, we observe $E(k)\sim k^{-2}$ and $E_{\rm c}(k)\sim k^{-2}$ for a decade or more.
-In the irrotationally-forced runs, $E_{\rm c}(k)\gg E_{\rm s}(k)$ across whole of the inertial range – the difference decreases as Ma increases.
-$E_{\rm s}(k)$, however, does not appear to have a well-defined power-law regime although it does appear to have a small scaling range with $E_{\rm s}(k)\sim k^{-3}$, marked by a dashed line.
-.
-B.2 Equal–time structure functions
-We calculate equal-time longitudinal structure functions of the total velocity, $S_{\rm p}(r)$, and its solenoidal, $S^{\rm s}_{\rm p}(r)$, and dilational, $S^{\rm c}_{\rm p}(r)$, see Fig.
-8.
-In each image, the regime of power-law fitting has been demarcated by the dashed lines.
-The scaling regimes of $S_{\rm p}(r)$, $S^{\rm c}_{\rm p}(r)$ and $S^{\rm s}_{\rm p}(r)$ are not exactly identical – there is partial overlap between them and all of them belong to the regime, $\eta\ll r\ll L_{\rm I}$.
-The fact that, even at a fixed Reynolds number, the scaling range of different quantities are different is found in incompressbile turbulence (Yakhot & Sreenivasan, 2004; Schumacher et al., 2007) and has also been observed in compressible turbulence Wang et al.
+Thus we obtain $M$ independent estimates for each exponent.
+The mean of these $M$ values is our estimates of the equal time exponets and the standard deviation is the error.
+For the scaling exponents obtained from the exit times, we introduce $2n$ tracer pairs, where $n=4096^{2}$, in the statistically stationary state and select only thoese whose initial and their final separation are within the inertial range.
+Thence we calculate the negative moment of the exit times and plot them, in log-log scale, as a function of $R$ and use linear fit to estimate the dynamic scaling exponents.
+We repeat this exercise for $M=20$ independent initial realization of the tracer pairs.
+For each exponent, its final estimate is its ensemble mean, and its error bars are defined by the standard deviation.
+These are the values plotted in Fig.
 (10).
-We select the range of scales over which we look for scaling of exit times as the range over which $S_{\rm p}(r)$ shows scaling.
-To calculate the scaling exponent $\zeta_{p}$ of $S_{\rm p}(r)$ we proceed the following way.
-We first consider the data from one single snapshot.
-Identify a scaling range by eye – this range is typically about a decade, but varies from one scheme of forcing to another.
-Then fit a straight line to the log-log plot of the structure functions to extract $\zeta_{p}$.
-We then collect many snapshots into an ensemble.
-The value of $\zeta_{p}$ we plot in Fig.
-(9) is the mean over this ensemble and the error is the standard deviation over the same ensemble.
-In statistical learning theory, this method is called BAGGing or Bootstrap AGGregation (Breiman, 1996; Mehta et al., 2019).
-The same method is applied to extract $\zeta_{p}^{\rm c}$ and $\zeta_{p}^{\rm s}$.
-Figure 9 shows that: (a) The solenoidal velocity fluctuations in S1 are less intermittent than those in S2; (b) At large $p$, $\zeta_{p}$ is much closer to $\zeta_{p}^{\rm c}$ because $S_{\rm p}(r)$ is dominated by the contribution from the longitudinal velocity differences across shocks which are predominantly irrotational; and (c) The saturation of $\zeta_{p}$ and $\zeta_{p}^{\rm c}$ at $p\gtrsim 2$ is brought about by the shocks.
-Figure 8: Equal-time structure functions.
-Log-log plots of structure functions: $S^{\rm c}_{\rm p}(r)$ (left column), $S^{\rm s}_{\rm p}(r)$ (middle column) and $S_{\rm p}(r)$ (right column) versus $r$ for $p=2$ (blue), $p=4$ (red) and $p=6$ (violet) in runs (a) S1, (b) S2, (c) C1 and (d) C2.
-The vertical dashed lines demarcate the region of power-law fitting in each case across which we extract $\zeta_{p}^{\rm c}$, $\zeta_{p}^{\rm s}$ and $\zeta_{p}^{\rm s}$.
-In each figure, the data has been averaged over a single non-equilibrium steady state (NESS) snapshot.
-Figure 9: Equal-time exponents.
-Scaling exponents, $\zeta_{p}$, $\zeta_{p}^{\rm s}$ and $\zeta_{p}^{\rm c}$ of the structure functions, $S_{\rm p}(r)$, $S^{\rm s}_{\rm p}(r)$ and $S^{\rm c}_{\rm p}(r)$, of different orders, $p$.
-Appendix C Lagrangian averages and dynamic exponents
-Our data for the moments of exit–times are averaged over $2n$ uniformly distributed pair separations for a single value of the inital separation $R$, with $n=4096^{2}$.
-Then the calculations are repeated for $25$ values of $R$ within the inertial range.
-The log-log plot of the moments of exit times as a function of $R$ are shown in Fig.
-(10).
-We run the DNS long enough (close to $3$ large–eddy–turnover–time) such that we have more than $10^{7}$ intervals that cross the thresholds for every value of $R$.
-This process is repeated for $M=10$ times by introducing the tracers to $M$ statistically independent initial conditions.
-Thus for every $R$ we have at least $10^{8}$ samples.
-We randomly divide our samples into $M=10$ ensembles and calculate the $M$ individual values of the exponents from each ensemble.
-The values of the exponents we quote are averaged over the ensemble and the the error we set is the standard deviation of the mean values obtained in each member.
-In statistical learning theory, this technique is known as BAGGing, or Bootstrap AGGregation (Breiman, 1996; Mehta et al., 2019).
-Therefore, from the sizes of the errorbars, we can conclude that the Lagrangian averages have converged.
-To check this further we calculate the exponents from $M/2$ randomly selected members of the ensemble and compare them with the exponents obtained from all ($M$) the members, see Fig.
-(11).
-The excellent overlap between the values of the dynamic exponents in these two cases demonstrate the convergence of the Lagrangian averages.
-Figure 10: Moments of Lagrangian exit times.
-Log-log plots of $\left\langle\tau_{\rm D}^{-p}(R)\right\rangle$ and $\left\langle\tau_{\rm H}^{-p}(R)\right\rangle$ versus $R$ for $p=2$ (blue), $p=4$ (red) and $p=6$ (violet) in runs (a) S1, (b) S2, (c) C1 and (d) C2.
-The vertical dashed lines demarkates the region of power–law fitting.
-Figure 11: Convergence of Lagrangian averages.
-The dynamic exponents calculated using half ($5$) the members of the ensemble, compared with the ones calculated using the whole ensemble ($M=10$), displayed in Fig.
-3 of the main text.
-The excellent overlap between the exponents and their errorbars, in every case affirms the convergence of the Lagrangian averages.
+This method of estimating parameters of a model and its errors from an ensemble based approach is often called BAGGing or Bootstrap AGGregation method, see e.g., Ref. [51] section 8.2.